### v2

# Spectral results for free random variables

Abstract.
Let $(\mathcal{A},\mathrm{tr})$ be a von Neumann algebra with a faithful, normal trace $\mathrm{tr}:\mathcal{A}\rightarrow\mathbb{C}.$ For each $a\in\mathcal{A},$ define
$S(\lambda,\varepsilon)=\mathrm{tr}[\log((a-\lambda)^{\ast}(a-\lambda)+\varepsilon)],\quad\lambda\in\mathbb{C},~\varepsilon>0,$
so that the limit as $\varepsilon\rightarrow 0^{+}$ of $S$ is the log potential of the Brown measure of $a.$ Suppose that for a fixed $\lambda\in\mathbb{C},$ the function
$\varepsilon\mapsto\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)=\mathrm{tr}[((a-\lambda)^{\ast}(a-\lambda)+\varepsilon)^{-1}]$
admits a real-analytic extension to a neighborhood of 0 in $\mathbb{R}.$ Then we will show that $\lambda$ is outside the spectrum of $a.$
We will apply this result to several examples involving circular and elliptic elements, as well as free multiplicative Brownian motions. In most cases, we will show that the spectrum of the relevant element $a$ coincides with the support of its Brown measure.
Hall’s research supported in part by a grant from the Simons Foundation
Ho’s research is supported in part by the NSTC grant 114-2115-M-001-005-MY3
1. Introduction
In this paper, we introduce a new characterization (Theorem 2.6) of the spectrum of an element in a tracial von Neumann algebra. We then apply this result to several examples, such as (1) the sum of an arbitrary element and a freely independent circular element, (2) more generally, the sum of an arbitrary element and a freely independent elliptic element, and (3) the product of a unitary element and a freely independent free multiplicative Brownian motion. Under suitable assumptions, we establish equality of the spectrum and the support of the Brown measure for these examples.
We now describe the origins of this line of research in the work of Leonard Gross. Let $K$ be a connected compact Lie group, where the unitary group $K=U(N)$ will be a key example. The paper [13] of Gross proved ergodicity for the action of the finite-energy loop group over $K$ on the continuous loop group with the pinned Wiener measure. A by-product of Gross’s proof was a Fock-space or “chaos” decomposition of the $L^{2}$ space over $K$ with respect to a heat kernel measure $\rho_{t}$. This result then motivated the introduction by Hall [16] of the Segal–Bargmann transform for $K.$ The transform is a unitary map from $L^{2}(K,\rho_{t})$ onto a holomorphic $L^{2}$ space of functions on the complexification $K_{\mathbb{C}}$ of $K.$ In the case $K=U(N),$ we have $K_{\mathbb{C}}=GL(N;\mathbb{C}),$ the group of all $N\times N$ invertible matrices over $\mathbb{C}.$
A paper of Gross and Malliavin [14] then gave a stochastic construction of the Segal–Bargmann transform over $K,$ using the Brownian motions in $K$ and $K_{\mathbb{C}}$ and methods from [13]. Finally, Biane [4] essentially took the construction of Gross and Malliavin for the case $K=U(N)$ and $K_{\mathbb{C}}=GL(N;\mathbb{C})$ and took the limit as $N\rightarrow\infty.$ Biane’s work indicated a close relationship between the free unitary Brownian motion (large-$N$ limit of Brownian motion in $U(N)$) and the free multiplicative Brownian motion (large-$N$ limit of Brownian motion in $GL(N;\mathbb{C})$).
Biane’s work then provided the motivation and technical tools for work of Hall–Kemp [23] computing the support of the Brown measure of the free multiplicative Brownian motion and then work of Driver–Hall–Kemp [9] computing the Brown measure itself. (Here “Brown measure” [6] is a von Neumann algebra construction that mimics the notion of eigenvalue distribution in random matrix theory.) The paper [9] introduced a new PDE method for computing Brown measures, which has then been used in subsequent works of Ho–Zhong [25], Hall–Ho [19, 20], Demni–Hamdi [8], and Eaknipitsari–Hall [10]. We also mention the work of Zhong [34], which does not use the PDE method but obtains similar formulas using free probability methods. The present paper also uses the PDE method in the applications of our general result.
We now discuss the results of the current paper. The Brown measure is defined for an element in a tracial von Neumann algebra $\mathcal{A},$ that is, a von Neumann algebra with a faithful, normal trace. (See Section 2.1 for details.) In general, the (closed) support of the Brown measure of $a$ is contained in the spectrum of $a.$ In many examples, the support of the Brown measure equals the spectrum and it is desirable to obtain conditions that would guarantee this equality. In the present paper, we introduce a new characterization (Theorem 2.6) of the spectrum of an element of a tracial von Neumann algebra, which is well suited for use with the PDE method of [9]. We then apply this result to get conditions on the spectrum in various examples.
We now briefly summarize the applications we will make of Theorem 2.6. We consider a element of the form $x+c$, where $c$ is circular and $x$ is freely independent of $x,$ or more generally $x+g,$ where $g$ is elliptic and freely independent of $x.$ Assume that the spectrum of $x$ coincides with the support of its Brown measure, which will happen, for example, if $x$ is normal. Then the spectrum of $x+c$ coincides with the support of its Brown measure (Corollary 4.8), and the same result holds more generally for $x+g$ (Theorem 4.15). Meanwhile, consider the general free multiplicative Brownian motion $b_{s,\tau}$ introduced in [20, Section 2.1] and let $u$ be a unitary element that is freely independent of $b_{s,\tau}.$ Then the spectrum of $ub_{s,\tau}$ coincides with the support of its Brown measure (Theorem 5.8). In particular, we determine the spectrum of the free multiplicative Brownian motion $b_{t},$ in the original form introduced by Biane [4, Section 4.2.1]. The spectrum of $b_{t}$ is equal to the support of its Brown measure, which is the closure of the domain $\Sigma_{t}$ in [9]. See Figure 1.
Figure 1. The region $\Sigma_{t}$ from [9] for $t=4$ (left), and a detail thereof (right).
2. A general result characterizing the spectrum
2.1. Brown measure
Let $(\mathcal{A},\mathrm{tr})$ be a tracial von Neumann algebra, that is a von Neumann algebra $\mathcal{A}$ together with a faithful, normal, tracial state $\mathrm{tr}:\mathcal{A}\rightarrow\mathbb{C}.$ Here “faithful” means that $\mathrm{tr}[a^{\ast}a]>0$ for all nonzero $a\in\mathcal{A},$ “normal” means that $\mathrm{tr}$ is continuous with respect to the weak operator topology, and “tracial” means that $\mathrm{tr}[ab]=\mathrm{tr}[ba]$ for all $a,b\in\mathcal{A}.$ For $a\in\mathcal{A},$ we let $\left|a\right|$ be the non-negative square root of $a^{\ast}a.$
If $a\in\mathcal{A}$ is a normal operator, we can define the law (or spectral distribution) $\mu_{a}$ of $a$ using the spectral theorem as
$\mu_{a}(E)=\mathrm{tr}[\nu_{a}(E)]$ (2.1)
for each Borel set $E,$ where $\nu_{a}$ is the projection-valued measure associated to $a$ by the spectral theorem (e.g., [17, Theorem 7.12]). The measure $\mu_{a}$ is the unique compactly supported probability measure on $\mathbb{C}$ satisfying
$\int_{\mathbb{C}}\lambda^{j}\bar{\lambda}^{k}~d\mu_{a}(\lambda)=\mathrm{tr}[a^{j}(a^{\ast})^{k}]$ (2.2)
for all non-negative integers $j$ and $k.$
Brown [6], extended the notion of law to elements that are not necessarily normal, as follows. For $\lambda\in\mathbb{C},$ we define
$s(\lambda)=\mathrm{tr}[\log(\left|a-\lambda\right|^{2})],$ (2.3)
which may be computed in terms of the law of $\left|a-\lambda\right|^{2}$ as
$s(\lambda)=\int_{0}^{\infty}\log(x)~d\mu_{\left|a-\lambda\right|^{2}}(x).$ (2.4)
The value of $s$ is defined to be -$\infty$ if $\mu_{\left|a-\lambda\right|^{2}}$ has positive mass at 0; the value of $s$ may also be -$\infty$ even if $\mu_{\left|a-\lambda\right|^{2}}$ has no mass at 0.
Brown showed that $s(\lambda)$ is finite for Lebesgue-almost-every value of $\lambda$ and is a subharmonic function of $\lambda.$ He then defined the Brown measure $\mathrm{Br}_{a}$ as
$\mathrm{Br}_{a}=\frac{1}{4\pi}\Delta s,$ (2.5)
where $\Delta$ is the distributional Laplacian.
Proposition 2.1.
Properties of $\mathrm{Br}_{a}$ include:
  1. (1)
    $\mathrm{Br}_{a}$ is a probability measure supported on the spectrum of $a.$
  2. (2)
    The function $s$ is the log potential of $\mathrm{Br}_{a},$ that is, the convolution of $\mathrm{Br}_{a}$ with the function $\log(\left|z\right|^{2}).$
  3. (3)
    $\mathrm{Br}_{a}$ agrees with $\mu_{a}$ if $a$ is normal.
  4. (4)
    We have
    $\int_{\mathbb{C}}\lambda^{j}~d\mathrm{Br}_{a}(\lambda)=\mathrm{tr}[a^{j}]$
    for all non-negative integers $j.$
  5. (5)
    If $\mathcal{A}$ is the space of $N\times N$ matrices with complex entries and $\mathrm{tr}$ is the normalized matrix trace, $\mathrm{tr}[a]=\frac{1}{N}\sum_{j=1}^{N}a_{jj},$ then $\mathrm{Br}_{a}$ is the empirical eigenvalue distribution of $a,$ namely
    $\mathrm{Br}_{a}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\lambda_{j}},$
    where $\{\lambda_{1},\ldots,\lambda_{N}\}$ are the eigenvalues of $A.$
We emphasize, however, that the Brown measure does not, in general, satisfy (2.2).
2.2. The spectrum and the support of the Brown measure
For any probability measure $\mu$ on $\mathbb{C},$ the support of $\mu,$ denoted $\mathrm{supp}(\mu),$ is the smallest closed set of full measure. In the case of the Brown measure of an element $a,$ we refer to $\mathrm{supp}(\mathrm{Br}_{a})$ as the Brown support of $a.$ In light of Point 1 of Proposition 2.1, the Brown support of any $a$ is contained in the spectrum of $a.$ Although, in many cases, the Brown support and the spectrum are actually equal, this is not always the case. Thus, it is desirable to identify tools that can allow us to prove equality of the Brown support and the spectrum in certain cases.
The following example shows that the spectrum and Brown support can differ for $R$-diagonal elements, that is, elements having the form of a Haar unitary times a freely independent non-negative element.
Example 2.2 (Haagerup–Larsen).
Suppose $h$ is a non-negative self-adjoint element such that the spectrum of $h$ contains 0 but $h$ has an $L^{2}$ inverse, meaning that
$\mathrm{tr}\left[h^{-2}\right]:=\int_{0}^{\infty}\frac{1}{\xi^{2}}~d\mu_{h}(\xi)<\infty.$
Let $x=uh,$ where $u$ is a Haar unitary that is freely independent of $h.$ Then the spectrum of $x$ is the disk
$\sigma(x)=\left\{\left.\lambda\in\mathbb{C}\right|\left|\lambda\right|^{2}\leq\mathrm{tr}[h^{2}]\right\},$
but the support of the Brown measure of $x$ is the annulus
[…]
The preceding result is a combination of Theorem 4.4(i) and Proposition 4.6 in [15].
The following elementary result says that for normal operators the spectrum and the Brown support agree.
Proposition 2.3.
If $a\in\mathcal{A}$ is normal, the closed support of $\mathrm{Br}_{a}=\mu_{a}$ is equal to the spectrum of $a.$
Proof.
If $P$ is a nonzero self-adjoint projection, then $\mathrm{tr}[P]=\mathrm{tr}[P^{2}]>0,$ by the faithfulness of the trace. It follows that the law $\mu_{a}$ of $a$ has the same sets of measure zero as the projection-valued measure $\nu_{a}$ associated to $a$ by the spectral theorem. Thus, the closed support of $\mu_{a}$ is the same as the closed support of $\nu_{a}.$ Part of the spectral theorem states that $\nu_{a}$ is supported on the spectrum of $a$ (which is a closed set), showing that the closed support of $\nu_{a}$ is contained in the spectrum.
We now show that the closed support of $\nu_{a}$ contains the spectrum of $a.$ If not, there would be a point $\lambda$ that is in $\sigma(a)$ but outside the closed support of $\nu_{a},$ which means that $\nu_{a}(D_{\lambda}(\varepsilon))=0$ for some open disk centered at $\lambda$ with radius $\varepsilon.$ Then consider the bounded function $f$ given by
$f(x)=\left\{\begin{array}[c]{cc}0&x\in D_{\lambda}(\varepsilon)\\ \frac{1}{x-\lambda}&x\notin D_{\lambda}(\varepsilon)\end{array}\right.$.
Then $f(a)$ is a bounded normal operator. Meanwhile, let $g(x)=x-\lambda,$ so that $g(a)=a-\lambda I.$ Then by the multiplicativity of the functional calculus for bounded measurable functions (e.g. Theorem 7.7 in [17]), we have
$f(a)g(a)=g(a)f(a)=(fg)(a).$
But the function $fg$ equals $1$ for $\nu_{a}$-almost-every $x,$ so $(fg)(a)=I.$ This shows that $f(a)$ is a bounded inverse of $g(a)=a-\lambda.$
2.3. The regularized log potential and its derivative
It is convenient to introduce the regularized log potential $S$ of $a\in\mathcal{A}$ as
$S(\lambda,\varepsilon)=\mathrm{tr}[\log(\left|a-\lambda\right|^{2}+\varepsilon)],\quad\lambda\in\mathbb{C},~\varepsilon>0.$ (2.6)
(See Section 11.5 of the monograph [28] of Mingo and Speicher.) Then $S$ is a $C^{\infty}$ function of both $\lambda$ and $\varepsilon$ and is subharmonic [28, Equation (11.8)] as a function of $\lambda$ for fixed $\varepsilon.$ For $\lambda\in\mathbb{C}$ fixed, $S(\lambda,\varepsilon)$ decreases as $\varepsilon$ decreases, so that the limit as $\varepsilon\rightarrow 0^{+}$ exists, possibly equal to $\infty.$ After separating the log function into its positive and negative parts and applying monotone convergence, we find that
$s(\lambda)=\lim_{\varepsilon\rightarrow 0^{+}}S(\lambda,\varepsilon)$
for all $\lambda\in\mathbb{C}.$
Remark 2.4.
Sometimes, a different convention is used, in which $\varepsilon$ is replaced by $\varepsilon^{2}$ on the right-hand side of (2.3). In the main results below, it is extremely important to distinguish between the $\varepsilon$ regularization” and the $\varepsilon^{2}$ regularization.” See Remark 2.7.
The function $S$ is a regularization of the log potential $s$ of $\mathrm{Br}_{a},$ in the sense that $S$ is a smooth function that approximates $s$ for small $\varepsilon.$ It is important to note, however, that $S$ cannot be computed from $s$; to compute $S,$ one needs information about the element $a$ that cannot (in general) be computed just from the function $s.$ (Thus, for example, $S$ cannot be computed as the convolution of $s$ with some mollifier function.) In particular, $S$ is not determined by $\mathrm{Br}_{a}$; if it were, it would also be determined by $s$, which is the log potential of $s.$
Although the function $S$ was introduced as a convenient regularization of the log potential $s$ of $\mathrm{Br}_{a},$ it plays a more fundamental role in certain Brown measure calculations. Specifically, Driver–Hall–Kemp [9] consider the log potential $S(t,\lambda,\varepsilon)$ of Biane’s free multiplicative Brownian motion $b_{t}.$ Then [9] shows that $S$ satisfies a PDE in which $\varepsilon$ appears as one of the variables. (See Section 5.1.) One cannot simply set $\varepsilon=0$ in the PDE because derivatives with respect to $\varepsilon$ appear. Further works using a PDE for the regularized log potential include those of Ho–Zhong [25], Hall–Ho [19, 20], Demni–Hamdi [8], and Eaknipitsari–Hall [10]. See also the first author’s expository discussion of the PDE method [18].
We consider also $\partial S/\partial\varepsilon.$ We use the general formula for the derivative of the trace of a logarithm,
$\frac{d}{du}\mathrm{tr}[\log(a(u))]=\mathrm{tr}[a(u)^{-1}],$
whenever $a(\cdot)$ is a differentiable function with values in the space of positive elements of $\mathcal{A}.$ (See [6, Lemma 1.1] or [18, Equation (25)].) Using this result, we compute that
$\frac{\partial S}{\partial\varepsilon}=\mathrm{tr}[(\left|a-\lambda\right|^{2}+\varepsilon)^{-1}]=\int_{0}^{\infty}\frac{1}{\xi+\varepsilon}~d\mu_{\left|a-\lambda\right|^{2}}(\xi).$ (2.7)
We then consider the behavior of $\partial S/\partial\varepsilon$ when $\varepsilon$ tends to zero and try to understand what it tells us about the Brown measure. We first note that,
$\lim_{\varepsilon\rightarrow 0^{+}}\frac{\partial S}{\partial\varepsilon}=\mathrm{tr}[\left|a-\lambda\right|^{-2}]$ (2.8)
where we define the right-hand side of (2.8) as the $\varepsilon\rightarrow 0^{+}$ of the last expression in (2.7), namely (by monotone convergence)
$\mathrm{tr}[\left|a-\lambda\right|^{-2}]=\int_{0}^{\infty}\frac{1}{\xi}~d\mu_{\left|a-\lambda\right|^{2}}(\xi).$ (2.9)
The quantity $\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)$ will typically blow up as $\varepsilon\rightarrow 0^{+},$ when $\lambda$ is in the support of the Brown measure of $a.$ We may consider for example, [9], which computes the Brown measure of the free multiplicative Brownian motion $b_{t}.$ In that setting, $\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)$ blows up like $1/\sqrt{\varepsilon}$ for $\lambda$ the interior of the support of the Brown measure of $b_{t},$ by Proposition (5.6) and Equation (5.12) in [9].
We emphasize that (2.9) can have a finite value even if $a-\lambda$ fails to have a bounded inverse; it is enough for $a-\lambda$ to have a inverse in the noncommutative $L^{2}$ space of operators $b$ with $\mathrm{tr}[b^{\ast}b]<\infty.$ Thus, the condition that $\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)$ has a finite limit as $\varepsilon\rightarrow 0^{+}$ does not, by itself, guarantee that $\lambda$ is outside the spectrum of $a.$ On the other hand, the following result of Zhong says that failure of $\partial S/\partial\varepsilon$ to blow up near $\lambda_{0}$ indicates that $\lambda_{0}$ is outside the Brown support of $a.$
Theorem 2.5 (Zhong).
If
$\lim_{\varepsilon\rightarrow 0^{+}}\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)=\mathrm{tr}\left[\left|a-\lambda\right|^{-2}\right]$
is finite for all $\lambda$ in some neighborhood of $\lambda_{0},$ then $\lambda_{0}$ is outside the support of the Brown measure of $a.$
See [34, Theorem 4.6]. This result is a strengthening of a result of Hall–Kemp [23, Theorem 1.2], which requires finiteness (and local boundedness) of the quantity $\mathrm{tr}[\left|(a-\lambda)^{2}\right|^{-2}].$
2.4. The main result
Our main result is a characterization of points $\lambda$ outside the spectrum of $a$ as the points where $\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)$ extends analytically in $\varepsilon$ to a neighborhood of $\varepsilon=0.$
Theorem 2.6.
Fix an element $a$ in a tracial von Neumann algebra $(\mathcal{A},\mathrm{tr})$ and define $S$ by (2.6). Suppose that for a fixed $\lambda\in\mathbb{C},$ the function
$\varepsilon\mapsto\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon),\quad\varepsilon>0,$ (2.10)
admits a real-analytic extension from $\varepsilon\in(0,\infty)$ to $\varepsilon\in(-\delta,\infty)$ for some $\delta>0.$ Then $a-\lambda$ is invertible, meaning that $\lambda$ is outside the spectrum of $a.$
Conversely, if $\lambda$ is outside the spectrum of $a,$ the map in (2.10) admits a real-analytic extension to $(-\delta,\infty)$ for some $\delta>0.$
We abbreviate the condition in the theorem as $\partial S/\partial\varepsilon$ is analytic in $\varepsilon$ at $\varepsilon=0.$” Note that if, for $\lambda$ fixed, $S$ itself is analytic in $\varepsilon$ at $\varepsilon=0,$ so is $\partial S/\partial\varepsilon.$
In Section 3, we will give several examples where $\partial S/\partial\varepsilon$ can be computed using the PDE method, giving restrictions on the spectrum of $a.$ In many of these examples, we will show that the spectrum of $a$ equals its Brown support.
Remark 2.7.
Suppose instead of the function $S,$ we consider the function
$\tilde{S}(\lambda,\varepsilon)=S(\lambda,\varepsilon^{2}),$
as in [20]. Suppose, for a fixed $\lambda,$ we can show that $\tilde{S}(\lambda,\varepsilon)$ has a real-analytic extension from $\varepsilon\in(-\infty,0)$ to $\varepsilon\in(-\infty,\delta)$ for some $\delta>0$ and that this extension is an even function of $\varepsilon$ on $(-\delta,\delta).$ Then $S(\lambda,\varepsilon)=\tilde{S}(\lambda,\sqrt{\varepsilon})$ will have also have a real-analytic extension to a neighborhood of $\varepsilon=0.$
We emphasize, however, that the existence of a real-analytic extension of $\tilde{S}(\lambda,\varepsilon)$ from $\varepsilon\in(-\infty,0)$ to $\varepsilon\in(-\infty,\delta)$ does not—without the evenness assumption—guarantee that $\lambda$ is outside the spectrum of $a.$ Indeed, Theorem 6.4 in [9] shows that such a real-analytic extension of $\tilde{S}$ can exist even for $\lambda$ in the Brown support of $a.$
We need the following (presumably well-known) result for tracial von Neumann algebras.
Lemma 2.8.
For all $b\in(\mathcal{A},\tau),$ if $b^{\ast}b$ is invertible, $b$ is also invertible.
Of course, this result does not hold for general operators on a Hilbert space. If $b$ is an isometry that is not surjective, then $b^{\ast}b=1$ is invertible but $b$ itself is not invertible. The lemma says that such examples cannot occur in a tracial von Neumann algebra.
Proof.
We use the polar decomposition to write $b=up,$ where $p$ is a non-negative self-adjoint operator and $u$ is a partial isometry with the kernel of $u$ equal to the kernel of $p.$ Now, $b$ must be injective in order for $b^{\ast}b$ to be invertible, and therefore $p$ must also be injective. Thus, $\ker u=\ker p=\{0\}.$ We conclude that $u$ is actually an isometry: $u^{\ast}u=1.$
Meanwhile, it is known [26, 6.1.3 Proposition] that $u$ and $p$ must also belong to $\mathcal{A}.$ Then by the cyclic property of the trace,
$\mathrm{tr}[1-uu^{\ast}]=\mathrm{tr}[1-u^{\ast}u]=\mathrm{tr}[1-1]=0.$ (2.11)
But $uu^{\ast}$ is the orthogonal projection onto the range of $u$. If this range were not the whole Hilbert space, $1-uu^{\ast}$ would be a nonzero, non-negative operator and (2.11) would contradict the faithfulness of the trace. Thus, $u$ is unitary and therefore invertible. But $p=(b^{\ast}b)^{1/2}$ is also invertible, so we conclude that $b=up$ is invertible. ∎
Proof of Theorem 2.6.
We denote by $f$ the real-analytic extension of the function in (2.10), which is real analytic on $(-\delta,\infty)$ for some $\delta>0.$ Then $f$ has a holomorphic extension, also called $f,$ from $(-\delta^{\prime},\delta^{\prime})$ to an open disk $D_{0}(\delta^{\prime})$ of radius $\delta^{\prime}$ centered at 0, for some $\delta^{\prime}\leq\delta.$
Let $G$ denote the Cauchy transform of $\left|a-\lambda\right|^{2},$ defined as
$G(z)=\mathrm{tr}\left[(z-\left|a-\lambda\right|^{2})^{-1}\right]=\int_{0}^{\infty}\frac{1}{z-\xi}~d\mu_{\left|a-\lambda\right|^{2}}(\xi),$ (2.12)
which is a holomorphic function of $z\in\mathbb{C}\setminus[0,\infty).$ Note from (2.7) that
$G(z)=-\frac{\partial S}{\partial\varepsilon}(\lambda,-z)=-f(-z)$
for $z<0.$ It follows that $G(z)$ agrees with f(z)-f(-z) on the connected open set $D_{0}(\delta^{\prime})\setminus[0,\delta^{\prime}).$ Thus, the restriction of $G$ to $D_{0}(\delta^{\prime})\setminus[0,\delta^{\prime})$ has a holomorphic extension to $D_{0}(\delta^{\prime}).$
Now, since $\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)$ is real valued for $\varepsilon<0,$ its real-analytic extension $f$ is also real valued on $(-\delta,\infty).$ Thus,
$\lim_{y\rightarrow 0^{+}}\operatorname{Im}[G(x+iy)]=-\lim_{y\rightarrow 0^{+}}\operatorname{Im}[f(-x-iy)]=0$
for all $x\in(-\delta,\delta),$ where the limit is locally uniform in $x$ by the continuity of $f$ on $D_{0}(\delta^{\prime}).$ Thus, by the Stieltjes inversion formula, the measure $\mu_{\left|a-\lambda\right|^{2}}$ is zero on $(-\delta,\delta).$ Proposition 2.3 then tells us that the spectrum of $\left|a-\lambda\right|^{2}$ does not include 0.
In the opposite direction, if $\lambda$ is outside the spectrum of $a,$ then $a-\lambda$ is invertible, so that $\left|a-\lambda\right|^{2}=(a-\lambda)^{\ast}(a-\lambda)$ is also invertible. Then for all $\varepsilon\in\mathbb{R}$ with $\left|\varepsilon\right|<\delta:=\left\|\left|a-\lambda\right|^{-2}\right\|,$ the inverse of $\left|a-\lambda\right|^{2}+\varepsilon$ exists, with
$(\left|a-\lambda\right|^{2}+\varepsilon)^{-1}$ $=\left|a-\lambda\right|^{-2}(1+\varepsilon\left|a-\lambda\right|^{-2})$
$=\left|a-\lambda\right|^{-2}\sum_{k=0}^{\infty}(-1)^{k}\varepsilon^{k}\left|a-\lambda\right|^{-2k}.$
Applying the trace to this relation gives a real-analytic function on $(-\delta,\delta)$ that agrees with $\frac{\partial S}{\partial\varepsilon}(\lambda,\varepsilon)$ on $(0,\delta).$
3. Applications
In the remainder of the paper, we study several examples where Theorem 2.6 can be used to give information about the spectrum of certain elements. The examples are mostly ones in which the PDE method introduced by Driver–Hall–Kemp [9] is used, such as papers by Hall–Ho [19, 20], Ho [24], Ho–Zhong [25], Demni–Hamdi [8], and Eaknipitsari–Hall [10]. We also analyze the examples studied by Zhong [34], where Zhong uses free-probability techniques instead of the PDE method but gets formulas similar to what one obtains from the PDE method.
In most cases, we show that the spectrum equals the Brown support, showing that the PDE method is even more powerful than was previously recognized.
We divide the examples into two broad classes, which we refer as “additive” and “multiplicative.”
3.1. Additive case
A semicircular element $x_{t}$ of variance $t>0$ in a tracial von Neumann algebra is a self-adoint element whose law is the semicircular measure on $2\sqrt{t},2\sqrt{t}]$, i.e., the measure with density $\frac{1}{2\pi t}\sqrt{4t-x^{2}}$ on this interval. A circular element $c_{t}$ of variance $t$ is then an element of the form
$c_{t}=\frac{1}{\sqrt{2}}(x_{t}+iy_{t}),$ (3.1)
where $x_{t}$ and $y_{t}$ are freely independent elements of variance $t.$ The Brown measure of $c_{t}$ is the uniform probability measure on a disk of radius $\sqrt{t}.$
An elliptic element is then an element of the form
$g=e^{i\theta}(ax+iby),$ (3.2)
where $a,$ $b,$ and $\theta$ are real numbers, with $a$ and $b$ not both zero, and where $x$ and $y$ are freely independent semicircular elements of variance 1. The $\ast$-distribution of $g$ is determined by the positive real number $t$ given by
$t=\mathrm{tr}[g^{\ast}g]$ (3.3)
and the complex number $\gamma$ given by
$\gamma=\mathrm{tr}[g^{2}].$ (3.4)
Then $\gamma$ satisfies
$\left|\gamma\right|\leq t$ (3.5)
and any pair $t>0$ and $\gamma\in\mathbb{C}$ satisfying (3.5) arises for some choice of $a,$ $b,$ and $\theta.$ (See [20, Section 2.1] or [34, Section 2.4].) We use the notation $g_{t,\gamma}$ to denote such an element.
The case $\gamma=0$ corresponds to the case $a=b$ in (3.2), in which case $g_{t,\gamma}=g_{t,0}$ is a circular element of variance $t.$ The case in which $\gamma=t$ corresponds to $\theta=b=0$ in (3.2) and gives a semicircular element of variance $a^{2}.$ We refer to models involving elliptic elements as “additive,” since the sum of two freely independent elliptic elements is again elliptic. Specifically, if $g_{t_{1},\gamma_{1}}$ and $g_{t_{2},\gamma_{2}}$ are freely independent elliptic elements, then
$g_{t_{1},\gamma_{1}}+g_{t_{2},\gamma_{2}}\overset{d}{=}g_{t_{1}+t_{2},\gamma_{1}+\gamma_{2}},$ (3.6)
where $\overset{d}{=}$ denotes equality in $\ast$-distribution. (Compare (3.13) in the multiplicative case.)
Ho and Zhong [25, Section 3] computed the Brown measure of an element of the form $x+c_{t},$ where $x$ is self-adjoint and freely independent of $c_{t},$ building on earlier results of Biane and Lehner [5, Section 3]. See Figure 2. Zhong then computed the Brown measure of $x+c_{t},$ where $x$ is freely independent of $c_{t}$ but otherwise arbitrary. See Figure 3.
Figure 2. Density plot of the Brown measure of $x+c_{t}$, where $x$ is self-adjoint with $\mu_{x}=\frac{1}{2}(\delta_{-1}+\delta_{1}),$ with $t=1.$
Figure 3. Density plot of the Brown measure of $x+c_{t},$ where $x$ is unitary and $\mu_{x}$ is supported at the third roots of unity, with equal masses, and $t=1.$
Meanwhile, Hall and Ho [19] computed the Brown measure of an element of the form $x+iy,$ where $x$ is self-adjoint, $y$ is semicircular, and $x$ and $y$ are freely independent. Ho [24] then computed the Brown measure of $x+g_{t,\gamma},$ where $x$ is self-adjoint, $g_{t,\gamma}$ is elliptic with $\gamma\in\mathbb{R}$, and $x$ and $g_{t,\gamma}$ are freely independent. See Figure 4. Finally, Zhong [34] computed the Brown measure of $x+g_{t,\gamma}$ where $x$ is arbitrary, $g_{t,\gamma}$ is a general elliptic element, and $x$ and $g_{t,\gamma}$ are freely independent.
Figure 4. Density plot of the Brown measure of $x+g_{t,\gamma}$ with $x$ self-adjoint and $\mu_{x}=\frac{1}{2}(\delta_{-1}+\delta_{1}),$ with $t=1$ and $\gamma=-i/2.$
The papers of Ho–Zhong [25], Hall–Ho [19], and Ho [24] are based on the PDE method introduced in [9]. The paper of Zhong [34], by contrast, uses techniques of free probability and subordination functions. Nevertheless, the formulas obtained by Zhong are parallel to what one gets when using the PDE method.
Theorem 3.1.
Let $g_{t,\gamma}$ be an elliptic element with parameters $t$ and $\gamma$ as in (3.3) and (3.4) and let $x$ be a self-adjoint element that is freely independent of $g_{t,\gamma}.$ Then the spectrum and the Brown support of $x+g_{t,\gamma}$ coincide:
$\sigma(x+g_{t,\gamma})=\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}}).$
The preceding result does not hold if $x$ is a arbitrary element freely independent of $g_{t,\gamma},$ even in the case that $g_{t,\gamma}$ is circular, as the following example shows.
Example 3.2.
Suppose $h$ is a non-negative self-adjoint element such that (1) the spectrum of $h$ contains 0 and (2) $h$ has an $L^{2}$ inverse, meaning that
$\mathrm{tr}\left[h^{-2}\right]:=\int_{0}^{a}\frac{1}{\xi^{2}}~d\mu_{h}(\xi)<\infty.$
Let $x=uh,$ where $u$ is a Haar unitary that is freely independent of $h.$ Then by Example 2.2, the spectrum of $x$ is a proper subset of its Brown support. Furthermore, for all sufficiently small $t,$ the spectrum of $x+c_{t}$ is a proper subset of its Brown support.
The proof of this result is deferred to Section 3.3.
A natural assumption to impose on $x$ is that the desired result (equality of spectrum and Brown support) should hold for $x$ itself.
Theorem 3.3.
Let $g_{t,\gamma}$ be an elliptic element with parameters $t$ and $\gamma$ as in (3.3) and (3.4) and let $x$ be an element that is freely independent of $g_{t,\gamma}$ but not necessarily self-adjoint. Assume that $\sigma(x)=\mathrm{supp}(\mathrm{Br}_{x}),$ which will hold, for example, if $x$ is normal (Proposition 2.3). Then
$\sigma(x+g_{t,\gamma})=\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}}).$
For more detailed statements of the preceding results, see Corollary 4.8 in the circular case and Theorem 4.15 in the general elliptic case.
For general $x,$ we can still prove the desired result, provided that $t$ is large enough.
Theorem 3.4.
Let $g_{t,\gamma}$ be an elliptic element with parameters $t$ and $\gamma$ as in (3.3) and (3.4) and let $x$ be an element that is freely independent of $g_{t,\gamma}$ but not necessarily self-adjoint. Then for all sufficiently large $t>0,$ we have
$\sigma(x+g_{t,\gamma})=\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}})$
for all $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t.$
See Corollary 4.9.
Remark 3.5.
In Theorems 3.1, 3.3, and 3.4, we may take $\gamma=0,$ in which case $g_{t,\gamma}$ becomes a circular element of variance $t$ and we obtain
$\sigma(x+c_{t})=\mathrm{supp}(\mathrm{Br}_{x+c_{t}}),$
under the stated hypotheses.
Remark 3.6.
The proofs of the preceding theorems rely on the prior computation of the relevant Brown support as the closure of a certain domain. Our new contribution is to show that there is no spectrum outside the closure of the domain. Then since the Brown measure of any element is supported on its spectrum, we obtain equality of the spectrum and Brown support.
In the circular case, $\mathrm{supp}(\mathrm{Br}_{x+c_{t}})$ was computed—by Ho–Zhong [25, Section 3] when $x$ is self-adjoint and by Zhong [34] when $x$ is arbitrary—as the closure $\overline{\Sigma}_{t}$ of a certain domain $\Sigma_{t}.$ Then by results of [34], $\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}})$ is the image of $\mathrm{supp}(\mathrm{Br}_{x+c_{t}})$ under a certain explicit map $\Phi_{t,\gamma}.$ See Section 4.3 for more information.
Further information about the domain $\Sigma_{t}$ in the preceding remark was obtained by Erdős and Ji in [11].
The preceding theorems will be proven in Section 4, in the following stages. We will start by analyzing $x+c_{t}$ in the self-adjoint case and then extend the arguments to $x+c_{t}$ where $x$ is not self-adjoint. Finally, for general $x,$ we will connect the case $x+g_{t,\gamma}$ to the case $x+c_{t}.$
3.2. Multiplicative case
We begin by giving a nonrigorous motivation for the model will introduce. Using (3.6), we can see that, for any $k,$
$g_{t,\gamma}\overset{d}{=}\frac{g_{t,\gamma}^{1}}{\sqrt{k}}+\cdots+\frac{g_{t,\gamma}^{k}}{\sqrt{k}},$ (3.7)
where $g_{t,\gamma}^{1},\ldots,g_{t,\gamma}^{k}$ are freely independent copies of $g_{t,\gamma}.$ We then make a “multiplicative” model by exponentiating, but where in the noncommutative setting, the correct way to exponentiate is to exponentiate the terms on the right-hand side of (3.7) separately and then multiply the results. Thus, we may consider
$\exp\left\{\frac{ig_{t,\gamma}^{1}}{\sqrt{k}}\right\}\exp\left\{\frac{ig_{t,\gamma}^{2}}{\sqrt{k}}\right\}\ldots\exp\left\{\frac{ig_{t,\gamma}^{k}}{\sqrt{k}}\right\}.$ (3.8)
Here the factor $i$ in the exponent is just a convention, which will give a nicer match of the parameters between the additive and multiplicative cases. (Note that $ig_{t,\gamma}$ is again an elliptic element, with parameters $t$ and $\gamma.$)
For large $k,$ we may reasonably hope to approximate each exponential in (3.8) using the Taylor series of the exponential through the quadratic order, considering instead
$\left(1+\frac{ig_{t,\gamma}^{1}}{\sqrt{k}}-\frac{(g_{t,\gamma}^{1})^{2}}{2k}\right)\left(1+\frac{ig_{t,\gamma}^{2}}{\sqrt{k}}-\frac{(g_{t,\gamma}^{2})^{2}}{2k}\right)\cdots\left(1+\frac{ig_{t,\gamma}^{k}}{\sqrt{k}}-\frac{(g_{t,\gamma}^{k})^{2}}{2k}\right).$ (3.9)
Now, for large $k,$ the term involving $(g_{t,\gamma}^{j})^{2}$ will be smaller than the term involving $g_{t,\gamma}^{j}$, because it has $k$ rather than $\sqrt{k}$ in the denominator. Nevertheless, the $(g_{t,\gamma}^{j})^{2}$ term is not negligible compared to the $g_{t,\gamma}^{j}$ term, because the $g_{t,\gamma}^{j}$ term has mean zero, while the $(g_{t,\gamma}^{j})^{2}$ generally has nonzero mean. We expect, however, that we can replace $(g_{t,\gamma}^{j})^{2}$ by $\mathrm{tr}[(g_{t,\gamma}^{j})^{2}]=\gamma$ in the large-$k$ limit, giving another model that should have the same large-$k$ behavior:
$\left(1+\frac{ig_{t,\gamma}^{1}}{\sqrt{k}}-\frac{\gamma}{2k}\right)\left(1+\frac{ig_{t,\gamma}^{2}}{\sqrt{k}}-\frac{\gamma}{2k}\right)\cdots\left(1+\frac{ig_{t,\gamma}^{k}}{\sqrt{k}}-\frac{\gamma}{2k}\right).$ (3.10)
To motivate the change from (3.9) to (3.10), we can compute that for a natural random matrix approximation $g_{t,\gamma}^{N}$ to $g_{t,\gamma},$ we have
$\mathbb{E}\{(g_{t,\gamma}^{N})^{2}\}=\mathbb{E}\{\mathrm{tr}[(g_{t,\gamma,}^{N})^{2}]\}I,$
where, here, $\mathrm{tr}$ denotes the normalized trace of a matrix.
At the rigorous level, we may define an elliptic Brownian motion $w_{t,\gamma}(r)$ by replacing the semicircular elements $x$ and $y$ in (3.2) by semicircular Brownian motions $x_{r}$ and $y_{r},$ that is, continuous processes with freely independent, semicircular increments. Then we may consider a free stochastic differential equation (3.11) based on (3.10) as
$db_{t,\gamma}(r)=b_{t}\left(1+i~dw_{t,\gamma}(r)-\frac{\gamma}{2}~dr\right),\quad b_{t,\gamma}(0)=1,$ (3.11)
where the $dr$ term is an Itô correction. We then define the free multiplicative Brownian motion with parameters $t$ and $\gamma$ as $b_{t,\gamma}$ as the value of $b_{t,\gamma}(r)$ at $r=1$:
$b_{t,\gamma}=\left.b_{t,\gamma}(r)\right|_{r=1}.$ (3.12)
See Section 2.1 in [20] for more information, where the parameter $s$ in [20] corresponds to $t$ here, while the parameter $\tau$ in [20] corresponds to $t-\gamma$ here. We refer to $b_{t,\gamma}$ as a “multiplicative” model, since it satisfies the multiplicative counterpart of (3.6):
$b_{t_{1},\gamma_{1}}b_{t_{2},\gamma_{2}}\overset{d}{=}b_{t_{1}+t_{2},\gamma_{1}+\gamma_{2}.}$ (3.13)
See Theorem 4.3 in [20].
The expression (3.8) represents a Wong–Zakai approximation (as in [31] or [30]) to $b_{t,\gamma}(1)$, obtained by making a piecewise-linear approximation $w_{t,\gamma}^{(k)}$ to the Brownian motion $w_{t,\gamma}$ and then solving (3.11) with $w_{t,\gamma}$ replaced by $w_{t,\gamma}^{(k)}$ (but without the Itô correction term). The expression (3.10) then represents a more numerically tractable approximation to $b_{t,\gamma}(1).$
In the case $\gamma=0,$ the free multiplicative Brownian motion was introduced by Biane. See [4, Section 4.2], where what we are calling the free multiplicative Brownian motion (with $\gamma=0$) is denoted $\Lambda_{t}.$ We use a special notation for the $\gamma=0$ case:
$b_{t}=b_{t,0}.$ (3.14)
Meanwhile, the case $\gamma=t$ corresponds to Biane’s free unitary Brownian motion $u_{t}$, introduced in [2]:
$b_{t,t}=u_{t}.$
The general form of the free multiplicative Brownian motion was introduced by Hall and Ho in [20], where $s$ and $\tau$ in [20] corresponds to $t$ and $t-\gamma,$ respectively, here.
Hall and Kemp showed that the support of Brown measure of $b_{t}$ is contained in the closure of a certain set $\Sigma_{t}$, which was introduced by Biane in [4, Section 4.2.6]. Driver, Hall, and Kemp [9] then computed the Brown measure of $b_{t}$ and showed that its support is exactly $\overline{\Sigma}_{t}.$ Ho and Zhong [25, Section 4] extended the results of [9] to compute the Brown measure of $ub_{t},$ where $u$ is a unitary element that is freely independent of $b_{t}$. Finally, Hall and Ho computed the Brown measure of $ub_{t,\gamma}$ for general $t$ and $\gamma.$ See Figures 5 and 6.
Meanwhile, Demni and Hamdi [8] studied the unitary Brownian motion $u_{t}$ multiplied by a non-negative self-adjoint initial condition $x$ (freely independent of $u_{t}$). In the case that $x$ is a self-adjoint projection, they identified a natural domain $\Sigma_{t}$ and showed that the support of the Brown measure of $xu_{t}$ is contained in $\{0\}\cup\overline{\Sigma}_{t}.$ Eaknipitsari and Hall [10] then extended the results of [8] to the case of $xb_{t,\gamma},$ where $x$ is a non-negative self-adjoint element freely independent of $b_{t,\gamma}.$
We now obtain information about the spectrum of $ub_{t,\gamma}$ and $xb_{t,\gamma}.$
Figure 5. Density plot of Brown measure of $ub_{t},$ where $u$ is unitary with $\mu_{u}$ supported at the fourth roots of unity, with equal mass, for $t=1.$
Theorem 3.7.
Let $b_{t}$ be the free multiplicative Brownian motion in (3.14) and let $u$ be a unitary element that is freely independent of $b_{t}.$ Then for all $t>0,$ we have
$\sigma(ub_{t})=\mathrm{supp}(\mathrm{Br}_{ub_{t}}).$
More generally, for any $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t,$ we have
$\sigma(ub_{t,\gamma})=\mathrm{supp}(\mathrm{Br}_{ub_{t,\gamma}}).$
For more detailed statements, see Theorem 5.4 in the case of $ub_{t}$ and Theorem 5.8 in the case of $ub_{t,\gamma}.$ As in the additive case, the proof of the theorem relies on the prior computation of the Brown support, in [9] for the case of $b_{t}$ itself, in [25, Section 4] for $ub_{t},$ and in [20] for $ub_{t,\gamma}$.
The case of a non-negative initial condition $x$ [8, 10] is conceptually similar to the case of a unitary initial condition, but more algebraically complicated. The algebraic complications prevent a rigorous computation of the Brown measure of $xb_{t,\gamma}.$ But [10] shows that the support of the Brown measure of $xb_{t,\gamma}$ is contained in $\{0\}\cup D_{t,\gamma}$ for a certain closed set $\gamma.$ We then show that “most” points outside $\{0\}\cup D_{t,\gamma}$ are outside the spectrum of $xb_{t,\gamma}.$ Precise statements may be found in Section 5.3.
Figure 6. Density plot of Brown measure of $ub_{t,\gamma},$ where $u$ is unitary with $\mu_{u}$ supported at the fourth roots of unity, with equal mass, for $t=1$ and $\gamma=-i/2.$
3.3. Verification of Example 3.2
Let $x=uh$ be as in the example, in which case Example 2.2 applies to $x.$ We will show that Example 2.2 continues to apply to $x+c_{t},$ for sufficiently small $t.$ To do this, we will show that 0 is in the spectrum of $\left|x+c_{t}\right|$ but that $\left|x+c_{t}\right|$ has an $L^{2}$ inverse.
For a measure $\mu$ on $\mathbb{R},$ we let $\tilde{\mu}$ be the symmetrization of $\mu,$ that is, the average of $\mu$ and its push-forward under the map $\xi\mapsto-\xi.$ We first note that the symmetrization of $\mu_{|c_{t}|}$ is the semicircular law $\mathrm{sc}_{t}$ of variance $t$ [28, p. 174]. By Proposition 3.5 of [15], $x+c_{t}$ is $R$-diagonal and the law of $|x+c_{t}|$, determined by its symmetrization, is the free convolution
$\tilde{\mu}_{|x+c_{t}|}=\tilde{\mu}_{h}\boxplus\mathrm{sc}_{t}.$ (3.15)
Now, by our assumption on $h$ and Proposition 2.3, 0 is in the support of $\mu_{h}$ and therefore also in the support of $\tilde{\mu}_{h}.$ It then follows from Proposition 2.2 in [7] that 0 is in the support of $\tilde{\mu}_{h}\boxplus\mathrm{sc}_{t}$ for all $t>0.$ (Taking $x=0$ in the cited proposition, the symmetry of the measures involved means that the quantity $u$ in the proposition must be zero. Then since 0 is in the support of $\tilde{\mu}_{h},$ the proposition tells us that 0 is in the support of $\tilde{\mu}_{h}\boxplus\mathrm{sc}_{t}.$) Thus, by (3.15), 0 is in the support of $\tilde{\mu}_{|x+c_{t}|}$ and thus, also, in the support of $\mu_{|x+c_{t}|}.$ Using Proposition 2.3 again, we conclude that 0 is in the spectrum of $\left|x+c_{t}\right|.$
We now show that $\left|x+c_{t}\right|$ has an $L^{2}$ inverse, for sufficiently small $t,$ using results of Biane [3] computing measures of the form $\mu\boxplus\mathrm{sc}_{t}.$ Consider the function $v_{t}$ defined by
[…]
where $v_{t}$ is continuous by [3, Lemma 2]. Then
$v_{t}(x)$ $=0\Longleftrightarrow\int_{\mathbb{R}}\frac{1}{(x-\xi)^{2}}~d\tilde{\mu}_{h}(\xi)\leq\frac{1}{t}$
$\Longleftrightarrow t\leq\left\|h^{-1}\right\|_{2}^{-2}.$ (3.16)
We then define an open set $\Omega_{t}$ inside the upper half-plane as the set of $x+iy$ with $y>v_{t}(x)$. Then define a holomorphic function on the upper half-plane by
$H_{t}(z)=z+tG_{\tilde{\mu}_{h}}(z),$
where $G_{\mu}$ denotes the Cauchy transform of a measure $\mu.$ According to Lemma 4 of [3], $H_{t}$ maps $\Omega_{t}$ injectively onto the upper half-plane. Furthermore, by (3.15) and [3, Proposition 2], we have that
$G_{\tilde{\mu}_{|x+c_{t}|}}(H_{t}(z))=G_{\tilde{\mu}_{h}}(z)$ (3.17)
for all $z\in\Omega_{t}.$
Now, since $\tilde{\mu}_{h}$ is symmetric,
$G_{\tilde{\mu}_{h}}(i\varepsilon)=\frac{1}{2}\int_{\mathbb{R}}\left(\frac{1}{i\varepsilon-\xi}+\frac{1}{i\varepsilon+\xi}\right)~d\tilde{\mu}_{h}(\xi)=-i\varepsilon\int_{\mathbb{R}}\frac{1}{\xi^{2}+\varepsilon^{2}}d\tilde{\mu}_{h}(\xi),$ (3.18)
so that, by monotone convergence,
$\lim_{\varepsilon\rightarrow 0^{+}}\frac{G_{\tilde{\mu}_{h}}(i\varepsilon)}{-i\varepsilon}=\int_{\mathbb{R}}\frac{1}{\xi^{2}}d\tilde{\mu}_{h}(\xi)=\left\|x^{-1}\right\|_{2}^{2}.$
Using (3.18), we calculate that
$H_{t}(i\varepsilon)=i\varepsilon\left(1-t\int_{\mathbb{R}}\frac{1}{\xi^{2}+\varepsilon^{2}}d\tilde{\mu}_{h}(\xi)\right).$
Since $\int_{\mathbb{R}}\frac{1}{\xi^{2}}~d\tilde{\mu}_{h}(\xi)$ is finite by assumption, we see that $H_{t}(i\varepsilon)\rightarrow 0$ as $\varepsilon\rightarrow 0^{+}.$ Furthermore, for $t<\|x^{-1}\|_{2}^{-2}$, we see from (3.16) that $i\varepsilon$ is in $\Omega_{t}$ for all $\varepsilon>0.$ Thus, using (3.17) and letting $\varepsilon\rightarrow 0^{+},$ we obtain
$\|(x+c_{t})^{-1}\|_{2}^{2}$ $=\lim_{\varepsilon\rightarrow 0^{+}}\frac{G_{\tilde{\mu}_{|x+c_{t}|}}(H_{t}(i\varepsilon))}{-H_{t}(i\varepsilon)}$
$=\lim_{\varepsilon\rightarrow 0^{+}}\frac{G_{\tilde{\mu}_{h}}(i\varepsilon)}{-i\varepsilon}\frac{i\varepsilon}{H_{t}(i\varepsilon)}$
$=\lim_{\varepsilon\rightarrow 0^{+}}\frac{G_{\tilde{\mu}_{h}}(i\varepsilon)}{-i\varepsilon}\frac{1}{1-t\frac{G_{\tilde{\mu}_{h}}(i\varepsilon)}{-i\varepsilon}}$
$=\|x^{-1}\|_{2}^{2}\frac{1}{1-t\|x^{-1}\|_{2}^{2}}.$
This shows that the inner radius of the Brown support of $x+c_{t}$ is
$\|(x+c_{t})^{-1}\|_{2}^{-1}=\sqrt{\|x^{-1}\|_{2}^{-2}-t},$
which is positive for all $t<\|x^{-1}\|_{2}^{-2}$.
4. Additive case
In this section, we provide more detailed statements and proofs for the results stated in Section 3.1.
4.1. The self-adjoint plus circular case
Ho and Zhong [25, Section 3] compute the Brown measure of an element of the form $x+c_{t},$ where $x$ is self-adjoint, $c_{t}$ is circular of variance $t,$ and $x$ and $c_{t}$ are freely independent. We now introduce some of the PDE techniques from [25] that we will use to prove a result about the spectrum of $x+c_{t}.$ (See also [18] for a gentle introduction to the PDE method.) We consider the regularized log potential of $x+c_{t}$, as in (2.3), which we write as
$S(t,\lambda,\varepsilon)=\mathrm{tr}[\log(\left|x+c_{t}-\lambda\right|^{2}+\varepsilon)],\quad\lambda\in\mathbb{C},~\varepsilon>0.$ (4.1)
According to [25, Proposition 3.2], $S$ satisfies the PDE
$\frac{\partial S}{\partial t}=\varepsilon\left(\frac{\partial S}{\partial\varepsilon}\right)^{2}$ (4.2)
with the initial condition
$S(0,\lambda,\varepsilon)=\mathrm{tr}[\log(\left|x-\lambda\right|^{2}+\varepsilon)].$ (4.3)
Note that no derivatives with respect to $\lambda$ appear, so we really have a PDE in $\varepsilon$ and $t,$ with $\lambda$ entering as a parameter in the initial conditions.
The PDE (4.2) is a first-order, nonlinear PDE of Hamilton–Jacobi type. We now briefly recap the method of characteristics as it applies to this equation. See Section 3.3 of the book [12] of Evans and Section 5.1 of [9] for more information. We introduce a “Hamiltonian” function by replacing $\partial S/\partial\varepsilon$ on the right-hand side of (4.2) with a “momentum” variable $p_{\varepsilon},$ with an overall minus sign:
$H(\varepsilon,p_{\varepsilon})=-\varepsilon p_{\varepsilon}^{2}.$
We then consider Hamilton’s equations for this Hamiltonian, meaning that we look for curves $\varepsilon(t)$ and $p_{\varepsilon}(t)$ satisfying
$\frac{d\varepsilon}{dt}$ $=\frac{\partial H}{\partial p_{\varepsilon}}(\varepsilon(t),p_{\varepsilon}(t))=2\varepsilon(t)p_{\varepsilon}(t)\quad$ (4.4)
$\frac{dp_{\varepsilon}}{dt}$ $=-\frac{\partial H}{\partial\varepsilon}(\varepsilon(t),p_{\varepsilon}(t))=-p_{\varepsilon}(t)^{2}.$ (4.5)
The initial condition $\varepsilon_{0}$ for $\varepsilon(t)$ is an arbitrary positive number,
$\varepsilon(0)=\varepsilon_{0},$
while the initial condition $p_{\varepsilon,0}$ for $p_{\varepsilon}(t)$ is obtained from the idea that the momentum variable $p_{\varepsilon}$ corresponds to $\partial S/\partial\varepsilon$:
$p_{\varepsilon,0}=\frac{\partial S}{\partial\varepsilon}(0,\lambda,\varepsilon_{0})=\mathrm{tr}[(\left|x-\lambda\right|^{2}+\varepsilon_{0})^{-1}].$ (4.6)
A curve of the form $t\mapsto\varepsilon(t)$, for some choice of $\varepsilon_{0},$ is called a characteristic curve of the PDE (4.2) with the initial condition (4.3). We then have the first and second Hamilton–Jacobi formulas. These assert that that if a solution to (4.4)–(4.5), with initial momentum given by (4.6), exists with $\varepsilon(t)>0$ up to some time $t_{\ast},$ then for all $t<t_{\ast},$ we have
$S(t,\lambda,\varepsilon(t))$ $=S(0,\lambda,\varepsilon_{0})+tH(\varepsilon_{0},p_{\varepsilon,0})$ (4.7)
$\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon(t))$ $=p_{\varepsilon}(t).$ (4.8)
The initial condition (4.6) ensures that the second Hamilton–Jacobi formula (4.8) holds at $t=0.$ Since we are interested in $\partial S/\partial\varepsilon,$ the second Hamilton–Jacobi formula will be more useful to us than the first.
Now, we can solve (4.5) as a separable equation, then plug the result into (4.4). Then (4.4) becomes separable as well, and we obtain the explicit formulas
$\varepsilon(t)$ $=\varepsilon_{0}(1-tp_{\varepsilon,0})^{2}$ (4.9)
$p_{\varepsilon}(t)$ $=\frac{p_{\varepsilon,0}}{1-tp_{\varepsilon,0}}.$ (4.10)
Once (4.9) is established, (4.10) is equivalent to the statement that
$\sqrt{\varepsilon(t)}~p_{\varepsilon}(t)=\sqrt{\varepsilon_{0}}p_{\varepsilon,0}.$ (4.11)
Note that when $t$ approaches the time
$t_{\ast}(\lambda,\varepsilon_{0})=\frac{1}{p_{\varepsilon,0}}=\frac{1}{\mathrm{tr}[(\left|x-\lambda\right|^{2}+\varepsilon_{0})^{-1}]},$ (4.12)
the solution of the system will cease to exist, because $p_{\varepsilon}(t)$ will approach infinity. We call $t_{\ast}(\lambda,\varepsilon_{0})$ the lifetime of the solution (4.9)–(4.10) to Hamilton’s equations (4.4)–(4.5).
Now, our goal is to understand the behavior of $\partial S/\partial\varepsilon$ near $\varepsilon=0,$ for a fixed $\lambda,$ using the Hamilton–Jacobi formulas (4.7)–(4.8). We therefore want to see what choice of initial condition $\varepsilon_{0}$ (where the value of $p_{\varepsilon,0}$ is determined by $\varepsilon_{0}$ as in (4.6)) will cause $\varepsilon(t)$ to be close to zero. Now, if we simply let $\varepsilon_{0}$ approach zero in (4.9), then $\varepsilon(t)$ will also approach zero—provided that the lifetime $t_{\ast}(\lambda,\varepsilon_{0})$ is at least $t$ in the limit as $\varepsilon_{0}$ tends to zero. If on the other hand, the $\varepsilon_{0}\rightarrow 0$ limit of $t_{\ast}(\lambda,\varepsilon_{0})$ is less than $t,$ it does not make sense to apply the Hamilton–Jacobi formula at time $t$ with $\varepsilon_{0}$ close to 0.
The preceding discussion leads us to consider the limit of $t_{\ast}(\lambda,\varepsilon_{0})$ as $\varepsilon_{0}\rightarrow 0,$ as follows:
$T(\lambda)=\lim_{\varepsilon_{0}\rightarrow 0^{+}}t_{\ast}(\lambda,\varepsilon_{0})=\frac{1}{\mathrm{tr}[\left|x-\lambda\right|^{-2}]},$ (4.13)
where $\mathrm{tr}[\left|x-\lambda\right|^{-2}]$ is interpreted as in (2.9). Since $x$ is self-adjoint, we can also write
$\mathrm{tr}[\left|x-\lambda\right|^{-2}]=\int_{\mathbb{R}}\frac{1}{\left|\xi-\lambda\right|^{2}}~d\mu_{x}(\xi).$ (4.14)
The quantity $\mathrm{tr}[\left|x-\lambda\right|^{-2}]$ cannot be zero but will be infinite for certain values of $\lambda.$ Thus, $T$ cannot be infinite but is zero when $\mathrm{tr}[\left|x-\lambda\right|^{-2}]$ is infinite. We then introduce a domain $\Sigma_{t}$ as
$\Sigma_{t}=\left\{\left.\lambda\in\mathbb{C}\right|T(\lambda)<t\right\}.$ (4.15)
We anticipate that the strategy of letting $\varepsilon_{0}\rightarrow 0$ will work outside the closure of $\Sigma_{t}.$
We now quote three technical results that we will need; their proof is given at the end of this subsection.
Lemma 4.1.
The function $T$ is upper semicontinuous on $\mathbb{C}$ and therefore the set $\Sigma_{t}$ is open.
Recall that a real-valued function $f$ on a metric space is said to be upper semicontinuous if for all $x,$
$\limsup_{y\rightarrow x}f(y)\leq f(x).$
Lemma 4.2.
If $x$ is self-adjoint, then for all $t>0,$ the spectrum of $x$ is contained in $\overline{\Sigma}_{t}.$
Lemma 4.3.
If $x$ is self-adjoint, then for all $t>0$ and $\lambda$ outside of $\overline{\Sigma}_{t},$ we have $T(\lambda)>t.$
Ho and Zhong show that for $\lambda$ outside $\overline{\Sigma}_{t},$ we can let $\varepsilon_{0}\rightarrow 0^{+}$ in (4.7), with the result that $\varepsilon(t)\rightarrow 0$ as well, giving
$\lim_{\varepsilon\rightarrow 0^{+}}S(t,\lambda,\varepsilon)=S(0,\lambda,0)=\mathrm{tr}[\log(\left|x-\lambda\right|^{2})],$
where $\mathrm{tr}[\log(\left|x-\lambda\right|^{2})]$ is well defined and harmonic for $\lambda$ outside $\bar{\Sigma}_{t},$ by Lemma 4.2. Thus, the Brown measure is zero outside $\overline{\Sigma}_{t}.$ There is a different analysis in [25, Section 3.2.2] to actually compute the Brown measure, inside $\Sigma_{t}$, but this does not concern us here—except for the result [25, Theorem 3.13] that the support of $\mathrm{Br}_{x+c_{t}}$ is equal to (not just contained in) $\overline{\Sigma}_{t}.$
We now refine the preceding analysis to show that points $\lambda$ outside $\overline{\Sigma}_{t}$ are outside the spectrum of $x+c_{t}.$
Theorem 4.4.
Let $c_{t}$ be circular of variance $t$, let $x$ be self-adjoint and freely independent of $c_{t},$ and consider the function $S$ in (4.1). Then for each $\lambda$ outside of $\overline{\Sigma}_{t},$ the function $\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon)$ is analytic at $\varepsilon=0.$ Thus, by Theorem 2.6, the spectrum of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}.$ Since [25, Theorem 3.13] tells us that the support of $\mathrm{Br}_{x+c_{t}}$ is exactly $\overline{\Sigma}_{t}$ (and since the Brown measure of any element is always supported on its spectrum), we conclude that the spectrum of $x+c_{t}$ coincides with its Brown support.
Proof.
We apply the second Hamilton–Jacobi formula (4.8) and the formula (4.10) for $p_{\varepsilon}(t)$ to get
$\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon(t))=\frac{p_{\varepsilon,0}}{1-tp_{\varepsilon,0}},$ (4.16)
where $p_{\varepsilon,0}$ is computed as a function of $\lambda$ and $\varepsilon_{0}$ by (4.6). We now fix some $\lambda$ outside $\overline{\Sigma}_{t}.$ We will first show that the right-hand side of (4.16) makes sense even when $\varepsilon_{0}$ is slightly negative. Then we will invert the relationship between $\varepsilon_{0}$ and $\varepsilon=\varepsilon(t)$ near $\varepsilon_{0}=\varepsilon=0$ and plug the result into (4.16) to obtain the desired analytic extension of $\partial S/\partial\varepsilon.$
We now fix some $\lambda$ outside $\overline{\Sigma}_{t}.$ By Lemma 4.2, $\lambda$ is outside the spectrum of $x$ and therefore $\left|x-\lambda\right|^{2}$ is invertible. In that case,
$p_{\varepsilon,0}=\mathrm{tr}[(\left|x-\lambda\right|^{2}+\varepsilon_{0})^{-1}]$
is actually well defined even when $\varepsilon_{0}$ is slightly negative. Thus, the map $\varepsilon_{0}\mapsto\varepsilon(t)$ is well defined and analytic in a neighborhood of $\varepsilon_{0}=0.$ Let us use the notation
$\tilde{p}_{\varepsilon,0}=\lim_{\varepsilon_{0}\rightarrow 0^{+}}p_{\varepsilon,0}=\mathrm{tr}[\left|x-\lambda\right|^{-2}]=\frac{1}{T(\lambda)}.$
Now, $T(\lambda)>t$ by Lemma 4.3, which means that $\tilde{p}_{\varepsilon,0}<1/t,$ so that $1-t\tilde{p}_{\varepsilon,0}>0.$ Thus, $\varepsilon_{0}\mapsto p_{\varepsilon}(t)$ is also well defined beyond $\varepsilon_{0}=0.$ Also,
$\left.\frac{\partial\varepsilon(t)}{\partial\varepsilon_{0}}\right|_{\varepsilon_{0}=0}$ $=\left.\left[(1-tp_{\varepsilon,0})^{2}+2\varepsilon_{0}(1-tp_{\varepsilon,0})\right]\right|_{\varepsilon_{0}=0}$
$=(1-t\tilde{p}_{\varepsilon,0})^{2}$
$>0.$
Thus, by the inverse function theorem, the map $\varepsilon_{0}\mapsto\varepsilon(t)$ has an analytic inverse map $E_{t}$ defined near 0. We may therefore construct an analytic function $f$ defined on $(-\delta,\delta)$ by
$f(\varepsilon)=\left.\frac{p_{\varepsilon,0}}{1-tp_{\varepsilon,0}}\right|_{\varepsilon_{0}=E_{t}(\varepsilon)}.$
By (4.16), this function agrees with $\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon)$ for $\varepsilon\in(0,\delta),$ so that $f$ gives the desired analytic extension. ∎
We now supply the proof of Lemmas 4.1, 4.2 and 4.3.
Proof of Lemma 4.1.
The function $t_{\ast}(\lambda,\varepsilon_{0})$ in (4.12) is continuous in $\lambda$ for $\varepsilon_{0}>0$. As $\varepsilon_{0}$ decreases to 0, $t_{\ast}(\lambda,\varepsilon_{0})$ decreases to $T(\lambda).$ It then follows from an elementary result (e.g., [32, Theorem 15.84]) that $T$ is upper semicontinuous and therefore that $\Sigma_{t}$ is an open set. ∎
Proof of Lemma 4.2.
Since $x$ is self-adjoint and therefore normal, we can apply Proposition 2.3 to conclude that the spectrum of $x$ coincides with the support of the law $\mu_{x}$ of $x.$ It then follows from the first paragraph of the proof of Theorem 3.8 in [25] that $\mathrm{supp}(\mu_{x})$ is contained in $\overline{\Sigma}_{t}$ for every $t.$ We can give different proof of this last statement as follows. Since $x$ is self-adjoint, $\mathrm{tr}[\left|x-\lambda\right|^{-2}]$ can be computed as in (4.14). (Compare how this quantity would be computed for general $x$ in (2.9).) Then by Lemma 4.5 in [34], the right-hand side of (4.14) is infinite for $\mu_{x}$-almost-every $\lambda\in\mathbb{C}.$ It follows from the definition (4.13) of $T$ that $T(\lambda)=0$ for $\mu_{x}$-almost-every $\lambda.$ Thus, by the definition (4.15) of $\Sigma_{t}$, we have that $\mu_{x}$-almost-every $\lambda$ is in $\Sigma_{t}.$ That is, $\Sigma_{t}$ is a set of full measure for $\mu_{x}$ and $\overline{\Sigma}_{t}$ is then a closed set of full measure, which must contain $\mathrm{supp}(\mu_{x}).$
Proof of Lemma 4.3.
We will show that for all $\lambda$ outside the spectrum of $x,$ we have
$\Delta\left(\frac{1}{T(\lambda)}\right)=4\mathrm{tr}[\left|(x-\lambda)^{2}\right|^{-2}]>0.$ (4.17)
Then, in light of Lemma 4.2, (4.17) will hold for all $\lambda$ outside $\overline{\Sigma}_{t}.$ We will conclude that, outside $\overline{\Sigma}_{t},$ the function $1/T$ cannot have a weak local maximum and the function $T$ cannot have a weak local minimum. Now, for $\lambda$ outside $\overline{\Sigma}_{t},$ we certainly have $T(\lambda)\geq t.$ If $T(\lambda)$ were equal to $t,$ then at all nearby points $\lambda^{\prime},$ we would have $T(\lambda^{\prime})\geq t,$ or else $\lambda$ would be in the closure of $\Sigma_{t}.$ But then $\lambda$ would be a weak local minimum for $T,$ which we have shown to be impossible.
We now verify (4.17). Fix $\lambda$ outside the spectrum of $x$ and use (4.13) to write
$\frac{1}{T(\lambda)}=\mathrm{tr}[\left|x-\lambda\right|^{-2}]=\mathrm{tr}[(x-\lambda)^{-1}(x^{\ast}-\bar{\lambda})^{-1}].$
Now, by the standard formula for the derivative of the inverse (e.g., [18, Equation (24)]), we have
$\frac{\partial}{\partial\lambda}(x-\lambda)^{-1}$ $=-(x-\lambda)^{-1}\left(\frac{\partial}{\partial\lambda}(x-\lambda)\right)(x-\lambda)^{-1}=(x-\lambda)^{-2}$
$\frac{\partial}{\partial\bar{\lambda}}(x-\lambda)^{-1}$ $=-(x-\lambda)^{-1}\left(\frac{\partial}{\partial\bar{\lambda}}(x-\lambda)\right)(x-\lambda)^{-1}=0,$
with similar formulas for the derivative of $(x^{\ast}-\bar{\lambda})^{-1}.$ Thus, differentiating under the trace, we get
$\Delta\mathrm{tr}[(x-\lambda)^{-1}(x^{\ast}-\bar{\lambda})^{-1}]$ $=4\frac{\partial^{2}}{\partial\bar{\lambda}\partial\lambda}\mathrm{tr}[(x-\lambda)^{-1}(x^{\ast}-\bar{\lambda})^{-1}]$
$=\mathrm{tr}[(x-\lambda)^{-2}(x^{\ast}-\bar{\lambda})^{-2}],$
as claimed. ∎
4.2. Arbitrary plus circular
We now consider the circular case of Zhong’s paper, $x+c_{t}$, where $x$ is freely independent of $c_{t}$ but otherwise arbitrary. We consider the function $T$ and the domain $\Sigma_{t}$ as in (4.13) and (4.15), but where we no longer assume that $x$ is self-adjoint. Lemma 4.1 still holds, with the same proof. But in this generality, our methods do not allow us to prove Lemma 4.2—that the spectrum of $x$ is inside $\overline{\Sigma}_{t}$. Thus, the proof of Theorem 4.4 breaks down at this point. Indeed, the conclusion of Theorem 4.4 is false for general $x,$ as Example 3.2 shows. In the example, the spectrum of $x$ is, by Example 2.2, a disk. But for small $t,$ the closed domain $\overline{\Sigma}_{t}$ is an annulus, so that $\sigma(x)$ is not contained in $\overline{\Sigma}_{t}.$
What we can prove is the following.
Theorem 4.5.
Let $c_{t}$ be a circular element of variance $t,$ let $x$ be another element (not necessarily self-adjoint) that is freely independent of $c_{t}.$ For all $\lambda\in\mathbb{C},$ if (1) $\lambda$ is outside the spectrum of $x,$ and (2) $T(\lambda)>t,$ then $\lambda$ is outside the spectrum of $x+c_{t}.$
Proof.
Although the paper [25] assumes that the element $x$ is self-adjoint, the derivation of the PDE (4.2) does not use this assumption. We may therefore attempt to follow the argument in the previous subsection. Now, the function $T$ in (3.3) is continuous outside the spectrum $\sigma(x)$ of $x.$ Thus, if $\lambda$ is outside $\sigma(x)$ and satisfies $T(\lambda)>t,$ then $T$ cannot be in the closure of the set $\Sigma_{t}=\{T<t\}$. Thus, under the assumptions of the theorem, the point $\lambda$ is outside $\overline{\Sigma}_{t},$ is outside $\sigma(x),$ and satisfies $T(\lambda)>t.$ At this point, the proof of Theorem 4.4 goes through without change. ∎
We now investigate how we can apply Theorem 4.5. In this investigation, the following result will be useful.
Lemma 4.6.
If $\lambda$ is outside the spectrum of $x$ and outside $\overline{\Sigma}_{t},$ then $T(\lambda)>t.$
Proof.
If we assume that $\lambda$ is outside the spectrum of $x,$ then the proof of Lemma 4.3 (from the self-adjoint case) goes through without change. ∎
For our first application of Theorem 4.5, we simply make the conclusion of Lemma 4.2 (from the case that $x$ is self-adjoint) an assumption.
Corollary 4.7.
Let $c_{t}$ be a circular element of variance $t,$ let $x$ be another element (not necessarily self-adjoint) that is freely independent of $c_{t}.$ Assume that, for some fixed $t,$ the spectrum of $x$ is contained in $\overline{\Sigma}_{t}.$ Then the spectrum of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}.$
Since [34, Theorem B] tells us that the support of $\mathrm{Br}_{x+c_{t}}$ is exactly $\overline{\Sigma}_{t}$ (and since the Brown measure of any element is always supported on its spectrum), we conclude that
$\sigma(x+c_{t})=\mathrm{supp}(\mathrm{Br}_{x+c_{t}})=\overline{\Sigma}_{t}.$
Proof.
If $\sigma(x)\subset\overline{\Sigma}_{t},$ then by Lemma 4.6, Theorem 4.5 will apply to every point outside $\overline{\Sigma}_{t}.$
Since our goal is ultimately to prove that the spectrum and Brown support of $x+c_{t}$ are equal, it is natural to assume that this condition holds at $t=0,$ that is, that the spectrum and Brown support of $x$ are equal.
Corollary 4.8.
If the spectrum and Brown support of $x$ coincide, then the spectrum of $x$ is contained in $\overline{\Sigma}_{t}$ and Corollary 4.7 tells us that
$\sigma(x+c_{t})=\mathrm{supp}(\mathrm{Br}_{x+c_{t}})=\overline{\Sigma}_{t}.$
Proof.
By (4.13) and (4.15), we have $\mathrm{tr}[\left|x-\lambda\right|^{-2}]\leq 1/t$ for $\lambda$ outside $\overline{\Sigma}_{t}.$ Then, by the last part of Theorem 4.6 in [34], the Brown support of xx—which by assumption equals the spectrum of xx—is contained in $\overline{\Sigma}_{t}.$ Thus, Corollary 4.7 applies. ∎
Even if the Brown support of $x$ is a proper subset of the spectrum of $x,$ we will still have that $\sigma(x)$ is inside $\overline{\Sigma}_{t}$ for all sufficiently large $t.$
Corollary 4.9.
For a fixed $x,$ the condition $\sigma(x)\subset\overline{\Sigma}_{t}$ holds for all sufficiently large $t$ and thus for all sufficiently large $t,$ we have
$\sigma(x+c_{t})=\mathrm{supp}(\mathrm{Br}_{x+c_{t}})=\overline{\Sigma}_{t}.$
Proof.
The function $T$ is upper semicontinuous by Lemma 4.1. Thus, $T$ achieves a maximum $T_{\max}$ on the compact set $\sigma(x),$ by an elementary property of upper semicontinuous functions. Thus, $\sigma(x)$ is contained in $\Sigma_{t}$ (and therefore also in $\overline{\Sigma}_{t}$) for all $t>T_{\max}.$
Remark 4.10.
Although our proof of Theorem 4.5 (following the proof of Theorem 4.4) uses the PDE method, we could also alternatively use results from Zhong’s paper [34], which is based on subordination-function methods. For example, we may look at Equation (3.13) in [34]. We identify $w$ there with $\sqrt{\varepsilon_{0}}$ here and $\varepsilon$ there with $\sqrt{\varepsilon(t)}$ here. Then after rearranging slightly, the formula in [34] becomes
$\sqrt{\varepsilon(t)}=\sqrt{\varepsilon_{0}}\left(1-t\mathrm{tr}\left[(\left|x-\lambda\right|^{2}+\varepsilon_{0})^{-1}\right]\right),$
which agrees with our formula (4.9) for $\varepsilon(t).$ We then consider the $\gamma=0$ case of Eq. (3.23) in [34], in which case, the quantity $z$ there equals $\lambda.$ This equation then says, in our notation, that
$\sqrt{\varepsilon(t)}\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon(t))=\sqrt{\varepsilon_{0}}\frac{\partial S}{\partial\varepsilon}(0,\lambda,\varepsilon_{0}),$
which is equivalent to the second Hamilton–Jacobi formula (4.8) with $p_{\varepsilon}(t)$ described by (4.10) or (4.11).
4.3. Arbitrary plus elliptic
Zhong [34] considers an element of the form
$x+g_{t,\gamma}$
where $g_{t,\gamma}$ is as in Section 3.1 and $x$ is freely independent of $g_{t,\gamma}.$ The case $\gamma=0$ corresponds to the case $x+c_{t}$ discussed in the previous subsection.
Although this is not how Zhong attacks the problem, it is possible to analyze $x+g_{t,\gamma}$ using a PDE method, by adapting the results of [20] to the additive setting. We use the notation
$S(t,\gamma,\lambda,\varepsilon)$
for the regularized log potential (as in (2.3)) of the element $x+g_{t,\gamma}.$ The PDE for $S$ would then be
$\frac{\partial S}{\partial\gamma}=-\frac{1}{2}\left(\frac{\partial S}{\partial\lambda}\right)^{2},$ (4.18)
where $\partial/\partial\gamma$ and $\partial/\partial\lambda$ are the Cauchy–Riemann operators with respect to the complex variables $\gamma$ and $\lambda.$ If we take $\gamma$ to be a real number $u$ and take the real part of both sides of (4.18), we obtain a PDE in real-variable form:
$\frac{\partial S}{\partial u}=-\operatorname{Re}\left[\left(\frac{\partial S}{\partial\lambda}\right)^{2}\right]=-\frac{1}{4}\left[\left(\frac{\partial S}{\partial x}\right)^{2}-\left(\frac{\partial S}{\partial y}\right)^{2}\right],\quad\lambda=x+iy.$ (4.19)
(There is no real loss of generality in assuming $\gamma$ to be real, since we can multiply the $x+g_{t,\gamma}$ by a constant of absolute value 1 to eliminate the factor of $e^{i\theta}$ in (3.2), at which point $\gamma$ becomes real.) Note that no derivatives with respect to $t$ or $\varepsilon$ appear in the PDEs (4.18) and (4.19).
Remark 4.11.
The PDEs (4.18) and (4.19) also arise in the analysis of the evolution of roots of polynomials when the polynomials evolve according to the heat flow, as in [21] and [22].
Although there is a PDE that applies to the case of $x+g_{t,\gamma},$ Zhong instead uses methods of free probability and subordination functions. We now state the main result of Zhong about this case.
Theorem 4.12 (Zhong).
Fix $t>0$ and $\gamma\in\mathbb{C}$ satisfying (3.5) and define a map $\Phi_{t,\gamma}:\mathbb{C}\rightarrow\mathbb{C}$ by
$\Phi_{t,\gamma}(\lambda)=\lambda+\gamma G_{x+c_{t}}(\lambda),$ (4.20)
where $G_{x+c_{t}}$ is the Cauchy transform of the Brown measure of $x+c_{t}$:
$G_{x+c_{t}}(\lambda)=\int_{\mathbb{C}}\frac{1}{\lambda-z}~d\mathrm{Br}_{x+c_{t}}(z).$ (4.21)
Then $\Phi_{t,\gamma}$ is continuous and the Brown measure $\mathrm{Br}_{x+g_{t,\gamma}}$ of $x+g_{t,\gamma}$ is the push-forward of $\mathrm{Br}_{x+c_{t}}$ under $\Phi_{t,\gamma}$:
$\mathrm{Br}_{x+g_{t,\gamma}}=(\Phi_{t,\gamma})_{\ast}(\mathrm{Br}_{x+c_{t}}).$
See Theorems C and D in [34]. In the case $x=0,$ we have that $\mathrm{Br}_{c_{t}}$ is uniform on an a disk, $\mathrm{Br}_{g_{t,\gamma}}$ is uniform on an ellipse, and the restriction of $\Phi_{t,\gamma}$ to the support disk of $\mathrm{Br}_{c_{t}}$ is real linear. Theorem 4.12 says that if we fix the element $x$ and the parameter $t$ but vary the parameter $\gamma$ starting from $\gamma=0,$ the Brown measure of $x+g_{t,\gamma}$ varies in a nice way—as push-forward under the explicit map given in (4.20). This sort of push-forward behavior is sometimes referred to as the model deformation phenomenon: deforming the free probability model (in a specific way) deforms the Brown measure in computable fashion. The model deformation phenomenon was actually first observed by Hall and Ho [20] in the multiplicative setting; see Section 5.2.
We now state our first result about the spectrum of $x+g_{t,\gamma}.$
Proposition 4.13.
Fix $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t$ and let $\Phi_{t,\gamma}$ be as in (4.20). Assume that the spectrum $\sigma(x)$ of $x$ is contained in $\overline{\Sigma}_{t},$ which will hold if $x$ is normal or, more generally, if $\sigma(x)$ coincides with the Brown support of $x.$ Then for all $\lambda$ outside of $\overline{\Sigma}_{t},$ the point $\Phi_{t,\gamma}(\lambda)$ is outside the spectrum of $x+g_{t,\gamma}.$
Lemma 4.14.
Under the assumptions of the proposition, the map $\Phi_{t,\gamma}$ is injective on the complement of $\overline{\Sigma}_{t}$ and may be computed on $(\overline{\Sigma}_{t})^{c}$ as
$\Phi_{t,\gamma}(\lambda)=\lambda+\gamma G_{x}(\lambda),\quad\lambda\in(\overline{\Sigma}_{t})^{c}.$ (4.22)
Observe that (4.22) involves $G_{x}(\lambda),$ while (4.20) involves $G_{x+c_{t}}(\lambda).$ We note that in some cases, $\Phi_{t,\gamma}$ is actually a homeomorphism of the whole complex plane onto itself. This result holds, for example, in these two cases: (1) when $x$ is self-adjoint with $\left|\gamma\right|\leq t$ but $\gamma\neq t,$ and (2) when $\left|\gamma\right|<t$ and $x$ is $R$-diagonal. See Corollary 6.9 and Theorem 7.8 in [34]. On the other hand, if $x=0$ and $\gamma=t=1$, then $\overline{\Sigma}_{t}$ is the closed unit disk and the restriction of $\Phi_{t,\gamma}$ to this disk is the map $\lambda\mapsto 2\operatorname{Re}(\lambda),$ which is not injective. But even in this case, $\Phi_{t,\gamma}$ remains injective on $(\overline{\Sigma}_{t})^{c}$; it is the conformal map $\lambda\mapsto\lambda+1/\lambda$ from the complement of the closed unit disk to the complement of [2,2].[-2,2]. (Take $\gamma=1$ in Example 1.5 in [34].)
Proof of Lemma 4.14.
We first let $\varepsilon\rightarrow 0$ in the second part of Eq. (3.33) in [34], which tells us that $p_{\lambda}^{(0)}(w(\varepsilon))=p_{\lambda}^{c,(t)}(\varepsilon),$ where these quantities are defined in Notation 3.10 of [34]. Now, for $\lambda$ outside $\overline{\Sigma}_{t},$ the quantity $w(\varepsilon)$ will tend to zero as $\varepsilon\rightarrow 0$ by [34, Lemma 3.5]. But by our assumptions, $\lambda\in(\overline{\Sigma}_{t})^{c}$ is also outside the spectrum of $x.$ Thus, the quantity $p_{\lambda}^{(0)}(w(\varepsilon))$ in Notation 3.10 in [34] will converge to
$\mathrm{tr}[(\lambda-x)^{-1}]=G_{x}(\lambda).$
Meanwhile, by Lemma 5.11 in [34], the quantity $p_{\lambda}^{c,(t)}(\varepsilon)$ in Notation 3.10 tends to $G_{x+c_{t}}(\lambda)$ as $\varepsilon\rightarrow 0.$ Thus, letting $\varepsilon\rightarrow 0$ in Eq. (3.33) gives
$G_{x+c_{t}}(\lambda)=G_{x}(\lambda),\quad\lambda\in(\overline{\Sigma}_{t})^{c},$ (4.23)
and (4.22) follows.
To prove the claimed injectivity, we use an argument due to Zhong (personal communication), which he has kindly allowed us to reproduce here. The argument generalizes the proof of Lemmas 3 and 4 in [3], but some additional steps. Assume that $\Phi_{t,\gamma}(\lambda_{1})=\Phi_{t,\gamma}(\lambda_{2})$ for $\lambda_{1},\lambda_{2}\in(\overline{\Sigma}_{t})^{c}.$ Then, using (4.22), we have
$\lambda_{1}-\lambda_{2}=-\gamma(G_{x}(\lambda_{1})-G_{x}(\lambda_{2})),$
so that
$\left|\lambda_{1}-\lambda_{2}\right|=\left|\gamma\right|\left|G_{x}(\lambda_{1})-G_{x}(\lambda_{2})\right|.$ (4.24)
Now, using (4.21) with $x+c_{t}$ replaced by $x,$ we get
$G_{x}(\lambda_{1})-G_{x}(\lambda_{2})=\int_{\mathbb{C}}\frac{(\lambda_{2}-\lambda_{1})}{(\lambda_{1}-z)(\lambda_{2}-z)}~d\mathrm{Br}_{x}(z).$
Applying the Cauchy–Schwarz inequality then gives
$\left|G_{x}(\lambda_{1})-G_{x}(\lambda_{2})\right|\leq\left|\lambda_{1}-\lambda_{2}\right|\left(\int_{\mathbb{C}}\frac{1}{\left|\lambda_{1}-z\right|^{2}}~d\mathrm{Br}_{x}\int_{\mathbb{C}}\frac{1}{\left|\lambda_{2}-z\right|^{2}}~d\mathrm{Br}_{x}\right)^{1/2}.$ (4.25)
But according to [34, Theorem 4.6],
$\int_{\mathbb{C}}\frac{1}{\left|\lambda-z\right|^{2}}~d\mathrm{Br}_{x}\leq\mathrm{tr}\left[\left|x-\lambda\right|^{-2}\right].$
Furthermore, since we assume $\sigma(x)\subset\overline{\Sigma}_{t},$ the proof of Lemma 4.3 applies, showing that
$\mathrm{tr}\left[\left|x-\lambda\right|^{-2}\right]=\frac{1}{T(\lambda)}<\frac{1}{t}.$
for $\lambda$ outside $\overline{\Sigma}_{t}.$ Thus, (4.25) becomes
$\left|G_{x}(\lambda_{1})-G_{x}(\lambda_{2})\right|<\left|\lambda_{1}-\lambda_{2}\right|\frac{1}{t}$
and (4.24) becomes
$\left|\lambda_{1}-\lambda_{2}\right|<\left|\gamma\right|~\left|\lambda_{1}-\lambda_{2}\right|\frac{1}{t}.$ (4.26)
Since $\left|\gamma\right|\leq t,$ (4.26) would be a contradiction unless $\left|\lambda_{1}-\lambda_{2}\right|=0.$ Thus, we obtain the claimed injectivity. ∎
Proof of Proposition 4.13.
Fix $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t$. Consider the regularized log potential of $x+g_{t,\gamma}$:
$S(t,\gamma,\lambda,\varepsilon)=\mathrm{tr}[(\log(\left|x+g_{t,\gamma}-\lambda\right|^{2}+\varepsilon)].$
When $\gamma=0,$ we obtain the function $S(t,\lambda,\varepsilon)$ in the previous subsection. For $\varepsilon>0,$ define a regularized version of $\Phi_{t,\gamma},$ denoted $\Phi_{t,\gamma}^{(\varepsilon)}$ by
$\Phi_{t,\gamma}^{(\varepsilon)}(\lambda)=\lambda+\gamma G_{x+c_{t},\varepsilon}(\lambda).$
Here, $G_{x+c_{t},\varepsilon}$ is the Cauchy transform of the regularized Brown measure of $x+c_{t},$ obtained by taking the Laplacian of the regularized log potential $S(t,\lambda,\varepsilon)$ of $x+c_{t},$ as in (4.1). That is to say,
$G_{x+c_{t},\varepsilon}(\lambda)=\int_{\mathbb{C}}\frac{1}{\lambda-z}\frac{1}{4\pi}\Delta_{\lambda}S(t,\lambda,\varepsilon)~d^{2}z.$
Define
$z(t,\gamma,\lambda,\varepsilon)=\Phi_{t,\gamma}^{(\varepsilon)}(\lambda).$ (4.27)
Then we have
$\frac{\partial S}{\partial\varepsilon}(t,\gamma,z(t,\gamma,\lambda,\varepsilon),\varepsilon)=\frac{\partial S}{\partial\varepsilon}(t,0,\lambda,\varepsilon).$ (4.28)
This result is the first relation in [34, Corollary 3.11], where the relation between $z$ and $\lambda$ there is given by Equation (3.27) and the last displayed equation in the proof of Proposition 5.2. One may also obtain (4.28) from the Hamilton–Jacobi analysis of the PDE in (4.18) or (4.19). From that perspective, (4.28) amounts to second Hamilton–Jacobi formula (4.8), along with the fact that—since $\varepsilon$ does not appear explicitly on the right-hand side of (4.18) or (4.19)— $p_{\varepsilon}$ is a constant of motion.
Now, by Corollary 4.7, points $\lambda$ outside $\overline{\Sigma}_{t}$ are also outside the spectrum of $x+c_{t},$ from which it follows that $\left|x+c_{t}-\lambda\right|^{2}$ is invertible. Near any such $\lambda,$ the regularized log potential $S(t,\lambda,\varepsilon)$ of $x+c_{t}$ is defined and analytic in $\varepsilon,$ even for $\varepsilon$ slightly negative. We now define, for each fixed $t$ and $\gamma,$ a map $F$ given by
$F(\lambda,\varepsilon)=(z(t,\gamma,\lambda,\varepsilon),\varepsilon),$
where for $\lambda\in(\overline{\Sigma}_{t})^{c},$ we allow $\varepsilon$ to be slightly negative and where $z(t,\gamma,\lambda,\varepsilon)$ is as in (4.27).
We then consider the matrix of derivatives $F_{\ast}$ of $F$ at $\varepsilon=0$ with $\lambda$ outside $\overline{\Sigma}_{t},$ which will have the form
$F_{\ast}(\lambda,0)=\left(\begin{array}[c]{cc}(\Phi_{t,\gamma})_{\ast}&\ast\\ 0&1\end{array}\right).$
Now, the support of the Brown measure of $x+c_{t}$ is contained in $\overline{\Sigma}_{t}$ [25, Theorem 3.8], from which we can see that the map $\Phi_{t,\gamma}$ in (4.20) is holomorphic on $(\overline{\Sigma}_{t})^{c}.$ Thus, $(\Phi_{t,\gamma})_{\ast}$ is just the holomorphic derivative (a complex number), interpreted as a $2\times 2$ matrix. But by Lemma 4.14, $\Phi_{t,\gamma}$ is injective on $(\overline{\Sigma}_{t})^{c},$ which means that the holomorphic derivative can never vanish.
We conclude that $F_{\ast}(\lambda,0)$ is invertible for all $\lambda$ outside $\overline{\Sigma}_{t}.$ It follows that $F$ has a real-analytic inverse defined near $(\Phi_{t,\gamma}(\lambda),0).$ We can then use the $\lambda$-component of $F^{-1}(z,\varepsilon)$ to define a function $\Lambda(z,\varepsilon)$ such that $F(\Lambda(z,\varepsilon),\varepsilon)=(z,\varepsilon)$ for $(z,\varepsilon)$ in a neighborhood of $(\Phi_{t,\gamma}(\lambda),0).$ Then we have
$\frac{\partial S}{\partial\varepsilon}(t,\gamma,z,\varepsilon)=\frac{\partial S}{\partial\varepsilon}(t,0,\Lambda(z,\varepsilon),\varepsilon)$ (4.29)
for $\varepsilon>0$ and the right-hand side of (4.29) provides a real-analytic extension of $\frac{\partial S}{\partial\varepsilon}(t,\gamma,z,\varepsilon)$ to $\varepsilon$ in a neighborhood of 0. Therefore, Theorem 2.6 applies and $\Phi_{t,\gamma}(\lambda)$ will be outside the spectrum of $x+g_{t,\gamma}.$
Theorem 4.15.
Assume that the spectrum $\sigma(x)$ of $x$ is contained in $\overline{\Sigma}_{t},$ which will hold if $x$ is normal or, more generally, if $\sigma(x)$ coincides with the Brown support of $x.$ For all $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t,$ define
$E_{t,\gamma}=\Phi_{t,\gamma}(\overline{\Sigma}_{t}).$ (4.30)
Then the spectrum and the Brown support of $x+g_{t,\gamma}$ both agree with $E_{t,\gamma}.$ In particular, the spectrum and the Brown support of $x+g_{t,\gamma}$ are equal.
Lemma 4.16.
Suppose $\mu$ is a compactly supported probability measure on $\mathbb{C}$ and $\Phi:\mathbb{C}\rightarrow\mathbb{C}$ is a continuous map. Let $\Phi_{\ast}(\mu)$ denote the push-forward of $\mu$ under $\Phi.$ Then
$\mathrm{supp}(\Phi_{\ast}(\mu))=\Phi(\mathrm{supp}(\mu)).$
Proof.
In general, a point belongs to the support of a measure if and only if every neighborhood of the point has positive measure. Note that $\Phi(\mathrm{supp}(\mu))$ is compact and therefore closed. Thus, if $z$ is outside $\Phi(\mathrm{supp}(\mu)),$ some neighborhood $U$ of $z$ is disjoint from $\Phi(\mathrm{supp}(\mu)).$ Thus, $\Phi^{-1}(U)$ is an open set contained in the complement of $\mathrm{supp}(\mu),$ so that $\Phi^{-1}(U)$ has measure zero with respect to $\mu$ and $U$ has measure zero with respect to $\Phi_{\ast}(U).$ Thus, $z$ is not in $\mathrm{supp}(\Phi_{\ast}(\mu)).$
In the other direction, suppose $z$ is in $\Phi(\mathrm{supp}(\mu)),$ meaning that $z=\Phi(\lambda)$ for some $\lambda$ in $\mathrm{supp}(\mu).$ Then for every neighborhood $U$ of $z,$ the set $\Phi^{-1}(U)$ is open and contains $\lambda,$ so that $\mu(\Phi^{-1}(U))>0$ and, thus, $\Phi_{\ast}(\mu)(U)>0.$ Thus, $z$ is in $\mathrm{supp}(\Phi_{\ast}(\mu)).$
Lemma 4.17.
The map $\Phi_{t,\gamma}$ is continuous and maps $\mathbb{C}$ onto $\mathbb{C}.$
Proof.
The continuity of $\Phi_{t,\gamma}$ follows from Lemma 5.11 in [34]. We will then follow one of the standard proofs of the fundamental theorem of algebra, using the concept of the fundamental group (e.g., [29, Theorem 56.1]). The Cauchy transform $G_{x+c_{t}}(\lambda)$ in the definition (4.20) of $\Phi_{t,\gamma}(\lambda)$ behaves like $1/\lambda$ near infinity. Thus,
$\Phi_{t,\gamma}(\lambda)\approx\lambda+\frac{\gamma}{\lambda}\approx\lambda$ (4.31)
near infinity.
Suppose that some $z\in\mathbb{C}$ failed to be in the image of $\Phi_{t,\gamma}.$ Then $\Phi_{t,\gamma}$ would map $\mathbb{C}$ continuously into the punctured plane $\mathbb{C}\setminus\{z\}.$ Now, if we restrict $\Phi_{t,\gamma}$ to a large circle $C$ centered at the origin, then by (4.31), $\Phi_{t,\gamma}(C)$ will have winding number 1 around $z$ and will therefore be homotopically nontrivial in $\mathbb{C}\setminus\{z\}.$ But on the other hand, $\mathbb{C}$ is simply connected, so the image under $\Phi$ of any loop in $\mathbb{C}$ must be homotopically trivial in $\mathbb{C}\setminus\{z\}.$ We therefore have a contradiction. ∎
Proof of Theorem 4.15.
Zhong’s result in Theorem 4.12 says that $\mathrm{Br}_{x+g_{t,\gamma}}$ is the push-forward of $\mathrm{Br}_{x+c_{t}}$ under $\Phi_{t,\gamma}.$ Thus, Lemma 4.16 tells us that the support of $\mathrm{Br}_{x+g_{t,\gamma}}$ is the set $E_{t,\gamma}$ in (4.30). Now, every $z$ outside $E_{t,\gamma}$ has the form $\Phi_{t,\gamma}(\lambda)$ for some $\lambda\in\mathbb{C}$, by Lemma 4.17. But since $z$ is not in $E_{t,\gamma}=\Phi_{t,\gamma}(\overline{\Sigma}_{t}),$ this $\lambda$ cannot be in $\overline{\Sigma}_{t}.$ Thus, by Proposition 4.13, $z$ is outside the spectrum of $x+g_{t,\gamma}.$ We conclude that
$\sigma(x+g_{t,\gamma})\subset E_{t,\gamma}=\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}}).$
Since the reverse inclusion $\mathrm{supp}(\mathrm{Br}_{x+g_{t,\gamma}})\subset\sigma(x+g_{t,\gamma})$ is a general property of Brown measures, we have the desired equality. ∎
5. Multiplicative case
5.1. The case of $ub_{t}$
We begin by considering an element of the form $ub_{t},$ where $b_{t}$ is the free multiplicative Brownian motion defined in Section 3.2 with $\gamma=0,$ and where $u$ is a unitary element that is freely independent of $b_{t}.$ We let $\mu_{u}$ denote the law of $u,$ as in (2.1).
We then introduce the regularized log potential of $ub_{t},$ as in (2.3),
$S(t,\lambda,\varepsilon)=\mathrm{tr}[\log(\left|ub_{t}-\lambda\right|^{2}+\varepsilon)].$
According to [9, Theorem 2.7], $S$ satisfies the PDE
$\frac{\partial S}{\partial t}=\varepsilon\frac{\partial S}{\partial\varepsilon}\left(1+(\left|\lambda\right|^{2}-\varepsilon)\frac{\partial S}{d\varepsilon}-x\frac{\partial S}{\partial x}-y\frac{\partial S}{\partial y}\right),\quad\lambda=x+iy.$ (5.1)
(Although [9] assumes $u=1,$ the derivation of the PDE there does not use this assumption.)
The Hamilton–Jacobi analysis of the PDE (5.1) then proceeds similarly to the additive case in Sections 4.1 and 4.2. One important difference in the two cases is that $\lambda$ and derivatives of $S$ with respect to the real and imaginary parts of $\lambda$ now appear on the right-hand side of (5.1). We must then incorporate $\lambda$ and an associated momentum variable $p_{\lambda}$ into the Hamiltonian system, with the initial value of $p_{\lambda}$ given by
$p_{\lambda}(0)=\frac{\partial S}{\partial\lambda}(t,\lambda_{0},\varepsilon_{0})=-\mathrm{tr}[(ub_{t}-\lambda)^{\ast}((\left|ub_{t}-\lambda\right|^{2}+\varepsilon)^{-1}].$ (5.2)
Then the second Hamilton–Jacobi formula for $\partial S/\partial\varepsilon$ takes the form:
$\frac{\partial S}{\partial\varepsilon}(t,\lambda(t),\varepsilon(t))=p_{\varepsilon}(t).$ (5.3)
(Compare (4.8) in the additive case, where $\lambda$ does not depend on $t.$)
Now, according to Proposition 5.9 in [9], it is possible to solve for the function $p_{\varepsilon}(t)$ explicitly in terms of the initial conditions of the system. It is then possible to compute the $\varepsilon_{0}\rightarrow 0$ limit of the lifetime as
$T(\lambda)=\frac{1}{\tilde{p}_{\varepsilon,0}(\lambda)}\frac{\log(\left|\lambda\right|^{2})}{\left|\lambda\right|^{2}-1},$ (5.4)
where at $\left|\lambda\right|=1,$ we assign $\log(\left|\lambda\right|^{2})/(\left|\lambda\right|^{2}-1)$ its limiting value, namely 1. Here $\tilde{p}_{\varepsilon,0}$ is the initial value of the momentum $p_{\varepsilon},$ evaluated at $\varepsilon_{0}=0,$ namely
$\tilde{p}_{\varepsilon,0}(\lambda)=\mathrm{tr}\left[\left|\lambda-u\right|^{-2}\right]=\int_{S^{1}}\frac{1}{\left|\lambda-\xi\right|^{2}}~d\mu_{u}(\xi).$ (5.5)
These calculations do not depend on the assumption that $u=1$ in [9]. We then define
$\Sigma_{t}=\left\{\left.\lambda\in\mathbb{C}\right|~T(\lambda)<t\right\}.$
We now state three technical lemmas, parallel to the ones in Section 4.1, that we will use in the proof of our main result. Their proofs are given at the end of this subsection.
Lemma 5.1.
The function $T(\lambda)$ equals $\infty$ at $\lambda=0$ and is finite elsewhere. Furthermore, the function $T$ is upper semicontinuous on $\mathbb{C}$ and therefore the set $\Sigma_{t}$ is open.
Lemma 5.2.
For all $t>0,$ the spectrum of the unitary element $u$ is contained in $\overline{\Sigma}_{t}.$
Lemma 5.3.
For all $t>0$ and all $\lambda$ outside $\overline{\Sigma}_{t},$ we have $T(\lambda)>t.$
We now state the main result of this section.
Theorem 5.4.
For all $t>0,$ the spectrum of $ub_{t}$ is contained in $\overline{\Sigma}_{t}.$ Thus, since [25, Theorem 4.28] shows that $\mathrm{supp}(\mathrm{Br}_{ub_{t}})=\overline{\Sigma}_{t}$ and since the support of the Brown measure of any element is contained in its spectrum, we conclude that
$\sigma(ub_{t})=\mathrm{supp}(\mathrm{Br}_{ub_{t}}).$
Proof.
Since both $b_{t}$ [4, p. 265] and $u$ are invertible, 0 is not in the spectrum of $ub_{t}.$
Assume than that $\lambda$ is a nonzero point outside $\overline{\Sigma}_{t}.$ By Lemma 5.2, $\lambda$ is outside the spectrum of $u.$ In that case, the initial momentum $p_{\varepsilon,0}$ at the point $\lambda$ with $\varepsilon_{0}=0$ (as in (5.5)) is well defined and finite, and similarly for the initial momentum $p_{\lambda}$ in (5.2). Indeed, these initial conditions remain well defined even when $\varepsilon_{0}$ is slightly negative.
If we evaluate at $\varepsilon_{0}=0,$ the lifetime of the solution of the Hamiltonian system is $T(\lambda),$ which is greater than $t$ by Lemma 5.3. Then a general result about flows (e.g., the fact that the sets $M_{t}$ in [27, Theorem 9.12] are open) tells us that for $(\lambda_{0},\varepsilon_{0})$ in a neighborhood of $(\lambda,0),$ the lifetime of the solution with initial conditions $(\lambda_{0},\varepsilon_{0})$ will remain greater than t.t. [^1] This point is more subtle than it may appear because the formula for the lifetime of the Hamiltonian system in [9, Proposition 5.9] is only valid for $\varepsilon_{0}\geq 0.$ The difficulty is that when $\varepsilon_{0}<0,$ the blowup time of the whole system may be smaller than the blowup time in the formula for $p_{\varepsilon}(t)$. See Remark 5.10 in [9].
We may then consider a map $U_{t}$ given by
$U_{t}(\lambda_{0},\varepsilon_{0})=(\lambda(t),\varepsilon(t)),$
where the characteristic curves $\lambda(\cdot)$ and $\varepsilon(\cdot)$ are computed using the initial conditions $\lambda_{0}$ and $\varepsilon_{0}$ and where the map is defined and analytic in a neighborhood of $(\lambda,0).$ Then by the proof of Lemma 6.3 in [9], the Jacobian of $U_{t}$ at $(\lambda_{0},0)$ is invertible. Thus, $U_{t}$ has a real-analytic inverse defined near $(\lambda_{0},0).$
We then apply the second Hamilton–Jacobi formula
$\frac{\partial S}{\partial\varepsilon}(t,\lambda(t),\varepsilon(t))=p_{\varepsilon}(t;\lambda_{0},\varepsilon_{0})$
from [9, Equation (5.8)], where the notation means that $p_{\varepsilon}$ is computed using the initial conditions $\lambda_{0}$ and $\varepsilon_{0}.$ Then
$p_{\varepsilon}(t;U_{t}^{-1}(\lambda,\varepsilon))$
will be analytic in $\varepsilon$ in a neighborhood of $\varepsilon=0$ and will agree with $\frac{\partial S}{\partial\varepsilon}(t,\lambda,\varepsilon)$ when $\varepsilon>0.$ Thus, Theorem 2.6 applies and $\lambda$ is outside the spectrum of $ub_{t}.$
We conclude the argument by supplying the proofs of Lemmas 5.1, 5.2, and 5.3.
Proof of Lemma 5.1.
As we have noted, $T(\lambda)$ has a removable singularity at $\left|\lambda\right|=1.$ Furthermore, $p_{\varepsilon,0}(\lambda)$ can be infinity but cannot be zero. Thus, the only way $T(\lambda)$ can be infinite is when $\lambda=0.$ Meanwhile, according to Proposition 4.8 in [25], the function $T(\lambda)$ is the decreasing limit as $\varepsilon_{0}\rightarrow 0^{+}$ of a certain function $t_{\ast}(\lambda,\varepsilon_{0})$ and this function is continuous in $\lambda$ for each $\varepsilon_{0}.$ The claimed semicontinuity of $T$ then follows by an elementary result [32, Theorem 15.84] about semicontinuous functions. ∎
Proof of Lemma 5.2.
Since $u$ is unitary, Proposition 2.3 tells us that the spectrum of $u$ equals the support of the law $\mu_{u}$ of $u.$ Now, by Lemma 4.5 in [34], the quantity $p_{\varepsilon,0}(\lambda)$ in (5.5) is infinite for $\mu_{u}$-almost every $\lambda.$ Thus, $T(\lambda)=0$ for $\mu_{u}$-almost every, showing that $\Sigma_{t}$ is a set of full measure for $\mu_{u}.$ Thus, $\overline{\Sigma}_{t}$ is a closed set of full measure for $\mu_{u},$ showing that $\sigma(u)=\mathrm{supp}(\mu_{u})$ is contained in $\overline{\Sigma}_{t}.$
Proof of Lemma 5.3.
The claimed result is stated in Theorem 4.10 of [25]. There is, however, a small gap in the proof, concerning the case $\left|\lambda\right|=1,$ which we fill in here. Since $T(0)=\infty,$ we only consider nonzero $\lambda.$ Now, a point $\lambda$ with $T(\lambda)=t>0$ will be outside $\overline{\Sigma}_{t}$ if and only if $T$ has a weak local minimum at $\lambda$ (meaning that all $\lambda^{\prime}$ in a neighborhood of $\lambda$ have $T(\lambda^{\prime})\geq t,$ so that such points are outside $\Sigma_{t}$). According to [25, Lemma 4.15], the function $T(re^{i\theta}),$ with $\theta$ fixed, is strictly increasing for $1<r<\infty,$ and strictly decreasing for $0<r<1.$ Thus, any possible weak local minimum of $T$ would have to be at a point on the unit circle.
We now consider points on the unit circle that are outside $\overline{\Sigma}_{t}$ and thus (Lemma 5.2) outside $\sigma(u)=\mathrm{supp}(\mu_{u}).$ Putting $\left|\lambda\right|=1$ in (5.4), and following the proof of Proposition 3.5 in [33], we may then compute that
$\frac{d^{2}}{d\theta^{2}}\frac{1}{T(e^{i\theta})}$ $=\frac{d^{2}}{d\theta^{2}}\int_{S^{1}}\frac{1}{2(1-\cos(\theta-\phi))}~d\mu_{u}(e^{i\theta})$
$=\frac{1}{2}\int_{S^{1}}\frac{(2+\cos(\theta-\phi))}{(1-\cos(\theta-\phi))^{2}}~d\mu_{u}(e^{i\theta})$
$>0.$ (5.6)
Thus, $1/T$ cannot have a weak local maximum at $e^{i\theta}$ and $T$ cannot have a weak local minimum at $e^{i\theta}.$ (Note, however, that $T$ can have a weak local minimum on the unit circle, namely when it is zero—in which case, (5.6) becomes meaningless—but this cannot happen at points in the unit circle outside $\mathrm{supp}(\mu_{u}).$) ∎
5.2. The case of $ub_{t,\gamma}$
We consider the general free multiplicative Brownian motion $b_{t,\gamma}$ as defined in Section 3.2. We then consider the regularized log potential of $ub_{t,\gamma}$, as in (2.3),
$S(t,\gamma,\lambda,\varepsilon)=\mathrm{tr}[\log((ub_{t,\gamma}-\lambda)^{\ast}(ub_{t,\gamma}-\lambda)+\varepsilon)].$
We use results of Hall–Ho [20], where $s$ in [20] corresponds to $t$ here and where $\tau$ in [20] corresponds to $t-\gamma$ here. According to Theorem 4.2 of [20], the function $S$ satisfies the PDE
$\frac{\partial S}{\partial\gamma}=-\frac{1}{8}\left(1-\left(1-\frac{1}{2}\varepsilon\frac{\partial S}{\partial\varepsilon}-2\lambda\frac{\partial S}{\partial\lambda}\right)^{2}\right),$ (5.7)
where we have adjusted the PDE to the $\varepsilon$” regularization used here, rather than the $\varepsilon^{2}$” regularization used in [20]. Note that unlike the PDE (4.18) in the additive case, derivatives with respect to $\varepsilon$ appear on the right-hand side of (5.7).
We now introduce the multiplicative version of the push-forward map $\Phi_{t,\gamma}$ in Section 4.3. In the multiplicative setting, it is convenient to use the Herglotz function in place of the Cauchy transform. If $a$ is an element of a tracial von Neumann algebra, we define
$J_{a}(\lambda)=\frac{1}{2}\int_{\mathbb{C}}\frac{\xi+\lambda}{\xi-\lambda}~d\mathrm{Br}_{a}(\xi),$
whenever the integral converges. (The factor of $\frac{1}{2}$ is a convenient normalization that makes the formulas in the multiplicative case more similar to the ones in the additive case.) Note that $J_{a}$ is related to the Cauchy transform $G_{a}$ as
$J_{a}(\lambda)=\frac{1}{2}\int_{\mathbb{C}}\frac{\xi-\lambda+2\lambda}{\xi-\lambda}~d\mathrm{Br}_{a}(\xi)=\frac{1}{2}-\lambda G_{a}(\lambda).$ (5.8)
We then define a map $\Psi_{t,\gamma},$ analogous to the map $\Phi_{t,\gamma}$ in the additive case, by
$\Psi_{t,\gamma}(\lambda)=\lambda\exp\left\{\gamma J_{ub_{t}}(\lambda)\right\},$ (5.9)
where $b_{t}$ is the value of $b_{t,\gamma}$ at $\gamma=0.$ By the $\tau=s$ case of [20], this map agrees with the one denoted $\Phi_{s,\tau}$ in [20, Section 8]. The following result shows that the “model deformation phenomenon” holds in this setting. That is to say, as we vary $\gamma$ with $u$ and $t$ fixed, the Brown measure of $ub_{t,\gamma}$ changes in a very specific way, namely by push-forward under the map $\Psi_{t,\gamma}.$
Theorem 5.5 (Ho–Zhong, Hall–Ho).
For all $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t,$ the Brown measure of $ub_{t,\gamma}$ is the push-forward of $\mathrm{Br}_{ub_{t}}$ under the map $\Psi_{t,\gamma}.$
This result is due to Ho–Zhong [25, Corollary 4.30] in the case $\gamma=0$ and to Hall–Ho [20, Theorem 8.2] for $\gamma\neq 0.$
Proposition 5.6.
Fix $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t$ and let $\Psi_{t,\gamma}$ be as in (5.9). Then for all $\lambda$ outside of $\overline{\Sigma}_{t},$ the point $\Psi_{t,\gamma}(\lambda)$ is outside the spectrum of $ub_{t,\gamma}.$
We now give the multiplicative version of Lemma 4.14.
Lemma 5.7.
For all $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t,$ the map $\Psi_{t,\gamma}$ is injective on the complement of $\overline{\Sigma}_{t}$ and is given on this set by
$\Psi_{t,\gamma}(\lambda)=\lambda\exp\left\{\gamma J_{u}(\lambda)\right\},\quad\lambda\in(\overline{\Sigma}_{t})^{c}.$ (5.10)
Thus, $\Psi_{t,\gamma}$ coincides on $(\overline{\Sigma}_{t})^{c}$ with the holomorphic function denoted $f_{\gamma}$ in [20, Definition 2.2]. Furthermore, $\Psi_{t,\gamma}$ is defined and continuous on $\mathbb{C}.$
Note that (5.9) involves $J_{ub_{t}}$ but (5.10) involves $J_{u}.$
Proof.
The formula (5.10) follows from the $\tau=s$ case of [20, Theorem 6.1]. Once (5.10) is established, we see that $\Psi_{t,\gamma}$ coincides on $(\overline{\Sigma}_{t})^{c}$ with the function denoted $f_{s-\tau}$ in [20], where $s$ and $\tau$ in [20] correspond to $t$ and $t-\gamma$, respectively, here. Then the claimed injectivity follows from Theorem 3.8 in [20]. Note that this theorem assumes $\tau\neq 0,$ which means $\gamma\neq t$ in our current notation, but in light of Lemma 5.3, this assumption is only needed to ensure injectivity of $f_{s-\tau}$ on the boundary of $\Sigma_{t}$; injectivity on $(\overline{\Sigma}_{t})^{c}$ still holds when $\gamma=t.$
Continuity of $\Psi_{t,\gamma}=f_{\gamma}$ outside $\overline{\Sigma}_{t}$ follows from [20, Equation (3.8)]. This formula also allows computation of the limiting value as we approach the boundary of $\Sigma_{t}.$ Meanwhile, continuity of $\Psi_{t,\gamma}$ on $\overline{\Sigma}_{t}$ follows from the explicit formulas in [20, Proposition 8.3], which agrees by construction with the limiting value of $\Psi_{t,\gamma}=f_{\gamma}$ on the boundary. ∎
Proof of Proposition 5.6.
The proof follows the proof of Proposition 4.13 in the additive case. We use the Hamilton–Jacobi analysis for the PDE (5.7) as developed in [20, Section 5], but keeping in mind that $\varepsilon$ in [20] corresponds to $\sqrt{\varepsilon}$ here. Now, [20] uses $\gamma=t$ (i.e., $\tau=0$ in the notation of [20]) as the initial condition. But since we have already established Theorem 5.4 (corresponding to the case $\gamma=0$), it is convenient to use $\gamma=0$ as our initial condition.
We then have the second Hamilton–Jacobi formula from [20, Theorem 5.1] for the PDE (5.7), adapted to the notation used here:
$\frac{\partial S}{\partial\varepsilon}(t,\gamma,\lambda(\gamma),\varepsilon(\gamma))=p_{\varepsilon}(\gamma),$
where the curves $\lambda(\gamma),$ $\varepsilon(\gamma),$ and $p_{\varepsilon}(\gamma)$ are given explicitly in Eqs. (5.6)–(5.9) of [20]. Now, if we take $\gamma=0$ as our initial condition (and use the $\varepsilon$” regularization), then the formulas in [20, Equations (5.4) and (5.5)] for the initial momenta become
$p_{\lambda}(0)$ $=-\mathrm{tr}\left[(ub_{t}-\lambda_{0})^{\ast}(\left|ub_{t}-\lambda_{0}\right|^{2}+\varepsilon_{0})\right]$ (5.11)
$p_{\varepsilon}(0)$ $=\mathrm{tr}\left[(\left|ub_{t}-\lambda_{0}\right|^{2}+\varepsilon_{0})\right],$ (5.12)
where $\lambda_{0}$ and $\varepsilon_{0}$ are the initial values of $\lambda(\gamma)$ and $\varepsilon(\gamma)$, respectively.
We then appeal to Theorem 5.4, which says that the spectrum of $ub_{t}$ is contained in $\overline{\Sigma}_{t}.$ Then for $\lambda_{0}\in(\overline{\Sigma}_{t})^{c},$ the initial momenta in (5.11) and (5.12) remain well defined and finite even if $\varepsilon_{0}$ is slightly negative. Thus, Eqs. (5.6)–(5.9) of [20] make sense and depend analytically on $\lambda_{0}$ and $\varepsilon_{0},$ even for $\varepsilon_{0}$ slightly negative.
Now, if we take $\gamma=0$ as our initial condition, the formula for $\lambda(\gamma)$ at $\varepsilon_{0}=0$ in [20, Proposition 5.6] becomes
$\lambda(\gamma)=f_{\gamma}(\lambda_{0})=\Psi_{t,\gamma}(\lambda_{0}),$
where the second equality is from Lemma 5.7. Furthermore, at $\varepsilon_{0}=0,$ we have $\varepsilon(\gamma)=0,$ by [20, Proposition 5.6].
We then define a map $F$, for each fixed $t$ and $\gamma,$ by
$F(\lambda_{0},\varepsilon_{0})=(\lambda(\gamma),\varepsilon(\gamma)),$
where the curves $\lambda(\cdot)$ and $\varepsilon(\cdot)$ are computed using the initial conditions $\lambda_{0}$ and $\varepsilon_{0}.$ The Jacobian of this map at $\varepsilon_{0}=0$ has the form
$F_{\ast}(\lambda_{0},0)=\left(\begin{array}[c]{cc}(\Psi_{t,\gamma})_{\ast}&\ast\\ 0&\left.\frac{\partial\varepsilon(\gamma)}{\partial\varepsilon_{0}}\right|_{\varepsilon_{0}=0}\end{array}\right),$
where the quantity in the bottom right corner is easily seen to equal 1, using the explicit formula for $\varepsilon(\gamma)$ in [20, Equation (5.7)]. The rest of the argument proceeds as in the proof of Proposition 4.13, using Lemma 5.7 in place of Lemma 4.14. ∎
Theorem 5.8.
For all $t>0$ and $\gamma\in\mathbb{C}$ with $\left|\gamma\right|\leq t,$ the spectrum $\sigma(ub_{t,\gamma})$ of $ub_{t,\gamma}$ is the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ and $\sigma(ub_{t,\gamma})$ coincides with the support of the Brown measure of $ub_{t,\gamma}.$
When $\gamma\neq t,$ the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ is the set denoted $\overline{\Sigma}_{s,\tau}$ in [20, Definition 2.5], where $s$ and $\tau$ in [20] correspond to $t$ and $t-\gamma$, respectively, here. See also Figure 4 in [20]. When $\gamma=t$ (corresponding to $\tau=0$ in [20]), the image of $\overline{\Sigma}_{t}$ under $\Psi_{t,\gamma}$ is the support of the law of the unitary element $uu_{t},$ where $u_{t}$ is Biane’s free unitary Brownian motion. In the $\gamma=t$ case, the restriction of $\Psi_{t,\gamma}$ to $\overline{\Sigma}_{t}$ is the map written in [25, Corollary 4.30] as
$\lambda\mapsto\Phi_{t,\bar{\mu}}(r_{t}(\theta)e^{i\theta}),$
where $r_{t}(\theta)$ is the inner radius of $\overline{\Sigma}_{t}$ at angle $\theta=\arg\lambda.$
Proof of Theorem 5.8.
The proof is almost identical to proofs of Lemma 4.17 and Theorem 4.15 in the additive case. ∎
5.3. The case of $xb_{t,\gamma}$
Demni and Hamdi [8] considered an element of the form $pu_{t},$ where $u_{t}$ denotes Biane’s free unitary Brownian motion and where $p$ is nonzero self-adjoint projection, freely independent of $u_{t}.$ They showed that the support of the Brown measure of $pu_{t}$ is contained in $\{0\}\cup\overline{\Omega}_{t,\alpha}$ for a certain set $\Omega_{t,\alpha},$ which is bounded by a Jordan curve. Work of Eaknipitsari–Hall [10] generalizes this result to elements of the form $xb_{t,\gamma},$ where $x$ is a non-negative self-adjoint element, assumed to be nonzero and freely independent of $b_{t,\gamma}.$ The case in which $\gamma=t$ (so that $b_{t,\gamma}=u_{t}$) and $x$ is a projection corresponds to the results of [8].
The PDEs (5.1) and (5.7) still hold for the regularized log potentials of $xb_{t}$ and $xb_{t,\gamma},$ respectively; only the initial conditions change. We then define two different momenta
$\tilde{p}_{0}(\lambda)$ $=\mathrm{tr}[\left|x-\lambda\right|^{-2}]$
$\tilde{p}_{2}(\lambda)$ $=\mathrm{tr}[\left|x\right|^{2}\left|x-\lambda\right|^{-2}],$
where the tilde on $p$ indicates that we are computing with $\varepsilon_{0}=0.$ Then $\tilde{p}_{0}$ is the initial value of the momentum $p_{\varepsilon},$ in the limit as $\varepsilon_{0}\rightarrow 0$; compare (5.5) in the case of a unitary initial condition. Meanwhile $\tilde{p}_{2}$ is another similar function that arises in various computations. We then consider a function $T$ defined as
$T(\lambda)=\frac{\log\left(\frac{\left|\lambda\right|^{2}\tilde{p}_{0}(\lambda)}{\tilde{p}_{2}(\lambda)}\right)}{\left|\lambda\right|^{2}\tilde{p}_{0}(\lambda)-\tilde{p}_{2}(\lambda)},$
with a limiting value of $1/\tilde{p}_{2}(\lambda)$ when $\left|\lambda\right|^{2}\tilde{p}_{0}(\lambda)=\tilde{p}_{2}(\lambda).$ This function describes the lifetime of Hamiltonian system associated to the PDE (5.1), with initial condition given by the regularized log potential of the non-negative element $x,$ in the limit as $\varepsilon_{0}\rightarrow 0.$ See [10, Definition 3.6 and Proposition 3.7].
We note that if $x$ were unitary, then $\left|x\right|^{2}$ would equal $1$ and $\tilde{p}_{2}$ would simply equal $\tilde{p}_{0}.$ In that case, the formula for $T$ would simplify to
$T(\lambda)=\frac{1}{\tilde{p}_{2}(\lambda)}\frac{\log(\left|\lambda\right|^{2})}{\left|\lambda\right|^{2}-1},\quad(\text{if }\left|x\right|^{2}=1),$
which is just the formula for $T(\lambda)$ in the unitary case. This observation suggests (correctly!) that the case of a positive initial condition is much harder to analyze than the case of a unitary initial condition.
The function $T(\lambda)$ is initially defined when $\tilde{p}_{0}(\lambda)$ and $\tilde{p}_{2}(\lambda)$ are finite, which holds outside the spectrum of $x$ and, thus, outside $0,\infty)$. By the proof of Proposition 3.10(ii) in [10], if $\tilde{p}_{0}(r)=\infty$ for some $r\in(0,\infty),$ then
$\lim_{\theta\rightarrow 0}T(re^{i\theta})=0.$
We can, therefore, extend $T(\lambda)$ to be defined for all nonzero complex numbers $\lambda,$ with $T(\lambda)=0$ whenever $\tilde{p}_{0}(\lambda)=\infty.$
We then define a set $\Sigma_{t},$ similarly to the previous examples in this paper, as
$\Sigma_{t}=\left\{\left.\lambda\neq 0\in\mathbb{C}\right|T(\lambda)<t\right\}.$
Note that, by definition, 0 is not in $\Sigma_{t};$ the origin will always be analyzed as a special case.
We now record what is known about the Brown support of $xb_{t}$ and $xb_{t,\gamma}.$
  • Theorem 3.21 in [10] asserts that the support of the Brown measure of $xb_{t}$ is contained in $\{0\}\cup\overline{\Sigma}_{t}.$ Furthermore, if 0 is outside $\overline{\Sigma}_{t},$ then 0 is in the support of the Brown measure of $xb_{t}$ if and only if $\mu_{x}(\{0\})>0.$ On the other hand, since [10] does not compute the Brown measure itself, it is not known whether the Brown support of $xb_{t}$ fills up all of $\overline{\Sigma}_{t}.$
  • Proposition 4.12 in [10] asserts that the support of the Brown measure of $xb_{t,\gamma}$ is contained in $\{0\}\cup D_{t,\gamma}$ for a certain closed set $D_{t,\gamma}$ defined in (5.13) below, but it is not known whether the Brown support fills up all of $D_{t,\gamma}.$
We then consider the extent to which Lemmas 5.1, 5.2, and 5.3 hold in this setting.
  • Although the function $T$ is not known to be upper semicontinuous, the set $\Sigma_{t}$ is open [10, Proposition 3.17].
  • According to Corollary 3.13 of [10], $\sigma(x)\setminus\{0\}$ is contained in $\overline{\Sigma}_{t}.$
  • If $\lambda$ is a nonzero complex number outside of $\overline{\Sigma}_{t}$ and $\lambda$ is outside $(0,\infty),$ then $T(\lambda)>t$. (See the last part of Proposition 3.19 in [10].) However, we cannot exclude the possibility that there is some $\lambda$ outside $\overline{\Sigma}_{t}$ but in $(0,\infty)$ with $T(\lambda)=t.$ (Numerically, it appears that such points $\lambda$ do not exist.)
Since the precise Brown support of $xb_{t}$ or $xb_{t,\gamma}$ is not known, it is not possible to prove that the spectrum and the Brown support coincide. Furthermore, because of the possibility of points outside $\overline{\Sigma}_{t}$ with $T(\lambda)=t,$ we cannot exclude the possibility of spectrum outside $\{0\}\cup\overline{\Sigma}_{t}$ or $\{0\}\cup D_{t,\gamma}.$ We now state the results we are able to obtain
Theorem 5.9.
If $\lambda$ is a nonzero complex number outside $\overline{\Sigma}_{t}$ and $T(\lambda)>t,$ then $\lambda$ is outside the spectrum of $xb_{t}.$ Furthermore, if 0 is outside $\overline{\Sigma}_{t},$ then 0 is in the spectrum of $xb_{t}$ if and only if $\mu_{x}(\{0\})>0.$
Proof.
Assume 0 is outside $\overline{\Sigma}_{t}.$ If $\mu_{x}(\{0\})>0,$ then 0 is in the Brown support of $xb_{t}$ by [10, Theorem 3.21] and, therefore, 0 is in the spectrum of $xb_{t}.$ On the other hand, if $\mu_{x}(\{0\})=0$ (and 0 is outside $\overline{\Sigma}_{t}$) then by [10, Corollary 3.13], 0 is outside the spectrum of $x.$ Thus, $x$ is invertible. Since, also, $b_{t}$ is invertible [4, p. 265], $xb_{t}$ is invertible so that 0 is outside the spectrum of $xb_{t}.$
We then consider a nonzero $\lambda$ outside $\overline{\Sigma}_{t}$ and we assume $T(\lambda)>0.$ Then the proof of Theorem 5.4 applies without change and we conclude that $\lambda$ is outside the spectrum of $xb_{t}.$
Figure 7. The domains $\Sigma_{t}$ (left) and $D_{t,\gamma}$ (right) for $t=\gamma=2.$
We then define a function a holomorphic function $f_{\gamma}$ on the complement of $\overline{\Sigma}_{t}$ by
$f_{\gamma}(\lambda)=\lambda\exp\left\{\gamma J_{x}(\lambda)\right\},$
where $J_{x}$ is the Herglotz transform of $x$ as in (5.8). Outside $\overline{\Sigma}_{t},$ the function $f_{\gamma}$ plays the role of the map $\Psi_{t,\gamma}$ from the case of a unitary initial condition. (Compare Lemma 5.7.) According to [10, Proposition 4.14], $f_{\gamma}$ is injective on $(\overline{\Sigma}_{t})^{c}$ and $f_{\gamma}(\lambda)$ tends to infinity as $\lambda$ tends to infinity. The closed set $D_{t,\gamma}$ in [10, Proposition 4.12] is then defined by the condition
$(D_{t,\gamma})^{c}=f_{\gamma}((\overline{\Sigma}_{t})^{c}).$ (5.13)
See Figure 7.
Theorem 5.10.
Let $z$ be a nonzero complex number outside $D_{t,\gamma}$ and let $\lambda$ be the complex number outside $\overline{\Sigma}_{t}$ such that $f_{\gamma}(\lambda)=z.$ If $T(\lambda)>t,$ then $z$ is outside the spectrum of $xb_{t,\gamma}.$ Furthermore, if 0 is outside $\overline{\Sigma}_{t},$ then the point $0=f_{\gamma}(0)$ is in the spectrum of $xb_{t,}\gamma$ if and only if $\mu_{x}(\{0\})>0.$
Proof.
The analysis of 0 is similar to the proof of Theorem 5.9, using [10, Proposition 4.13] in place of [10, Theorem 3.21]. If $\lambda$ is a nonzero point outside $\overline{\Sigma}_{t}$ and we assume $T(\lambda)>0,$ then Theorem 5.9 tells us that $\lambda$ is outside the spectrum of $xb_{t}.$ Then the proof of Proposition 5.6 applies, using the injectivity of $f_{\gamma}$ obtained in [10, Proposition 4.14], showing that $f_{\gamma}(\lambda)$ is outside the spectrum of $xb_{t,\gamma}.$
Acknowledgments
We thank Jorge Garza-Vargas and Benson Au for useful discussions on a related problem that motivated this line of research. We thank Hari Bercovici for supplying the proof of Lemma 2.8 and we thank Ping Zhong for several helpful comments.